### v1

# Debate, Train, Evolve: Self-Evolution of Language Model Reasoning

Abstract
Large language models (LLMs) have improved significantly in their reasoning through extensive training on massive datasets. However, relying solely on additional data for improvement is becoming increasingly impractical, highlighting the need for models to autonomously enhance their reasoning without external supervision. In this paper, we propose Debate, Train, Evolve (DTE), a novel ground truth-free training framework that uses multi-agent debate traces to evolve a single language model. We also introduce a new prompting strategy Reflect-Critique-Refine, to improve debate quality by explicitly instructing agents to critique and refine their reasoning. Extensive evaluations on five reasoning benchmarks with six open-weight models show that our DTE framework achieve substantial improvements, with an average accuracy gain of 8.92% on the challenging GSM-PLUS dataset. Furthermore, we observe strong cross-domain generalization, with an average accuracy gain of 5.8% on all other benchmarks, suggesting that our method captures general reasoning capabilities.
1 Introduction
Over the past few years, the advancements in large language models (LLMs) have largely depended on training over massive datasets Abdin et al. (2024, 2025). However, eventually, we will approach a saturation point where feeding more data into these models may not further improve their reasoning capabilities Costello et al. (2025). This motivates a new research question: How can language models continue to improve without relying on additional external supervision?
Recent approaches attempt to overcome the data bottleneck by enabling models to generate and learn from synthetic data, which is generated by automatically expanding a small set of seed tasks into large synthetic instruction datasets Wang et al. (2022); Zeng et al. (2024). Other methods Madaan et al. (2023); Jiang et al. (2023); Gou et al. (2023); Peng et al. (2023); Zelikman et al. (2024); Costello et al. (2025) refine model-generated outputs through iterative self-feedback or preference optimization. Despite their effectiveness, these self-evolution strategies predominantly rely on judgments from a single model or a teacher-student configuration, often leading to confirmation bias and insufficient reasoning diversity.
To address these limitations, one promising direction emerged is multi-agent debate (MAD) Du et al. (2023). It involves multiple models independently generating and critically analyzing each other’s answers, helping to reveal subtle reasoning errors often overlooked by individual models Liang et al. (2023); Wang et al. (2024). Although MAD shows improved reasoning accuracy, current works predominantly use MAD as an inference-time technique Smit et al. (2023), requiring multiple models to be run simultaneously for each query. This substantially increases computational overhead and latency Subramaniam et al. (2025), making MAD impractical for large-scale deployments. This motivates our research question: Can we evolve a single model reasoning by fine-tuning on these debate traces?
Building upon this intuition, we propose Debate, Train, Evolve (DTE), a novel framework that combines the strengths of MAD with efficient single-model inference. Specifically, we introduce a ground-truth-free training approach in which a model learns from its own debate traces generated during MAD, thereby evolving autonomously over iterative training cycles. Our framework addresses key challenges of existing methods by extracting high-quality reasoning insights from diverse multi-agent interactions, thus avoiding single-model biases and computational inefficiencies.
First, we conduct a large-scale empirical analysis of MAD using open-source models, where we identify limitations of the original MAD prompting approach, particularly in smaller models Du et al. (2023). To address this, we propose a Reflect-Critique-Refine (RCR) prompting strategy, which explicitly forces agents to identify, critique, and correct reasoning errors in both their own and peers’ answers. Second, using this prompting strategy, we build our DTE framework (Figure 1). Finally, we find that models with parameters suffer accuracy loss Srivastava et al. (2025) after second evolution round; our controlled study shows that the problem correlates with large temperature-induced variance and high KL divergence from the base policy. Lowering the sampling temperature from 0.7 to 0.3 cuts the KL drift by and recovers up to 76% of the lost performance, preventing catastrophic forgetting in smaller models without extra supervision.
Our experiments show significant gains in reasoning performance across multiple datasets. Specifically, our evolved models show an average accuracy improvement of 8.92% on the challenging GSM-PLUS dataset compared to their original versions. Moreover, our framework achieves notable cross-domain generalization, enhancing model performance across datasets not seen during training. These results confirm that our Debate, Train, Evolve method successfully distills multi-agent debate’s insights into efficient single-model inference, bridging the gap between computational efficiency and advanced reasoning capabilities.
2 Related Work
Multi-Agent Debate Approaches
Du et al. (2023) first showed that letting several large models debate improves accuracy on maths, strategy, and factual QA without any new parameters. Later, Liang et al. (2023) highlighted the risk of degeneration-of-thought: a single agent quickly converges on one path, whereas a two-debater plus judge setup maintains diversity and outperforms GPT-4 on tricky arithmetic. RECONCILE (Chen et al., 2023) mixes agents from different model families, reaches consensus through confidence-weighted votes, and adds up to eleven points on seven reasoning benchmarks. Smit et al. (2023) shows that MAD beats sampling ensembles only after careful tuning. Finally, works like PREDICT (Park et al., 2024) apply multi-agent debate to tasks beyond QA, such as hate-speech classification, where agents reason under different guidelines. Recent advances further incorporate explicit reinforcement learning into the debate process. For example, the ACC-Collab framework Estornell et al. (2024) utilized an actor-critic approach to explicitly optimize agent collaboration, yielding superior performance on reasoning tasks.
Self-Evolution in Language Models
SELF-INSTRUCT (Wang et al., 2022) prompts GPT-3 to write 52000 novel instructions plus answers and then fine-tunes on its own output, reducing the gap to InstructGPT by thirty-three points on Super-Natural-Instructions without extra human labels. STaR (Zelikman et al., 2024) augments a few chain-of-thought exemplars by letting the model explain wrong answers in reverse, doubling CommonsenseQA accuracy for a 350M model. SELF-REFINE (Madaan et al., 2023) and the broader SELF framework (Lu et al., 2023) turn one model into writer, critic and re-writer, looping feedback at inference or during fine-tuning to improve on GSM8K by around seven points. Instruction-tuning variants refine the idea: Self-Refine Instruction-Tuning (Ranaldi and Freitas, 2024) pairs Llama-2 and Mistral students with large teacher rationales and then lets each student prefer its own better reasoning, closing the size gap on commonsense and math tasks. More recently, Think, Prune, Train, Improve (Costello et al., 2025) shows that careful filtering of self-generated traces can raise Gemma-2B to 58% on GSM8K and push Llama-3-70B beyond GPT-4o. These studies confirm that single-agent loops, with or without ground truth, can expand a model’s ability.
Despite these works, two things remain unexplored: 1) Fully autonomous, ground-truth-free self-evolution; 2) Integration of MAD into model evolution. Our work addresses this by the Debate, Train, Evolve framework, which combines MAD with self-supervised reinforcement learning (GRPO) to enable models to autonomously evolve their reasoning capabilities.
3 Method
Preliminary Study: Multi-Agent Debate (MAD)
First, we conduct a large-scale empirical study on MAD. Our initial objective was to understand how MAD affects performance across different reasoning tasks (eg., math, science), model families, and across scales, and to identify its limitations. Our ultimate goal was to generate high-quality reasoning traces using MAD. Following MAD Du et al. (2023) prompting strategy, however, our initial experiments show that the gains from MAD were minimal across reasoning benchmarks.
Our manual inspection uncovered two primary issues: (i) sycophancy: agents often discarded their correct answer and copied a other agent’s wrong but confident solution; (ii) verbosity bias: when two answers disagreed, agents preferred the longer rationale even if it was logically weak Saito et al. (2023). For example, on average, a three-agent Qwen-1.5B committee, sycophancy reached (28% of the debates have sycophancy rounds). These effects lead to a substantial fraction of [correct incorrect] transitions during debate.
3.1 Reflect–Critique–Refine Prompting
Inspired by previous works that had iterative self-feedback procedures Madaan et al. (2023); Gou et al. (2023); Peng et al. (2023), we propose Reflect-Critique-Refine (RCR) prompting, which mitigates this problem. Unlike the original MAD prompt that simply asks agents to revise their answer by looking at others, RCR prompting encourages agents to present confident reasoning. 1) Every agent must reflect by stating why its current answer could be wrong; 2) Each agent then critiques exactly two peer rationales, naming a specific flaw or a step it cannot verify; 3) Finally the agent may refine its answer—but only if it supplies at least one new reasoning step.
Phrases like “identify any errors” reliably trigger negative tokens (“error”, “mistake”, “step X is wrong”) which LLMs have learned during supervised finetuning. By specifying valid next moves (defend/correct/adopt), we implicitly shape the log-probability mass toward useful trajectories, shrinking the space of rambling answers. The single-step explanation requirement forces agents to think before copying, while the fixed quota of peer critiques limits needless verbosity. This prompting helped reduce sycophancy and improved performance.
Debate protocol
Let be a set of agents and a query. Round 0: each produces , an answer and a rationale. If consensus (all answers identical) occurs, the process stops. Otherwise we enter a debate round following Algorithm 1: agents receive all previous rationales, execute Reflect–Critique–Refine, and output . The debate ends when either (i) consensus is reached or (ii) a maximum of rounds is exhausted, after which the final answer is decided by simple majority vote.
3.2 Debate, Train, Evolve Framework
So far we have seen that RCR prompting yields high-quality debate traces and measurable accuracy gains. We now move from learning from these debate traces to train a single language model. This results in Debate, Train, Evolve (DTE), a closed loop that distills the collective insight of multiple agents into a single, faster model.
Framework overview
Let be a frozen base policy with parameters . At iteration we run Algorithm 1 on a batch of queries , producing a set of debate traces where is the original query, the consensus answer and a compressed rationale extracted from the highest-scoring agent. Extraction step keeps any explanation steps that appear in at least two agents or that introduce a symbolic manipulation absent from the previous round.
We then fine-tune a student policy on with Group Relative Policy Optimization Shao et al. (2024) and then place back into the agent pool, replacing its earlier version. The loop repeats until the mean validation reward stagnates for two iterations or a maximum of five cycles is reached. Algorithm 2 gives the exact procedure.
Reward shaping and GRPO
For a query and a student output we define the shaped reward
where is the indicator function, checks the <reasoning> <answer> XML template, and is the token length.
Weights are set to and .
GRPO maximises the clipped surrogate
where , , and . The first term is identical to PPO but uses the debate-derived reward; the second anchors the student to the base policy, a key factor for avoiding catastrophic forgetting Luo et al. (2025); Kotha et al. (2024). Gradients are estimated with the standard REINFORCE algorithm and a value network baseline.
| Model | GSM8K | GSM-Plus | ARC-Challenge | ||||||
| Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | |
| Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | |
| Qwen-2.5-1.5B | 62.77 | 72.33 | 73.09 (+10.32 ) | 42.00 | 53.33 | 55.92 (+13.92 ) | 69.21 | 68.52 | 68.36 (-0.85 ) |
| Qwen-2.5-3B | 84.08 | 85.14 | 86.05 (+1.97 ) | 61.75 | 68.00 | 69.50 (+7.75 ) | 83.53 | 84.64 | 83.95 (-0.42 ) |
| Qwen-2.5-7B | 90.67 | 91.21 | 88.32 (-2.35 ) | 68.62 | 74.17 | 74.71 (+6.09 ) | 87.22 | 91.64 | 90.89 (+3.67 ) |
| Qwen-2.5-14B | 92.80 | 93.33 | 93.74 (+0.94 ) | 71.79 | 77.25 | 78.88 (+7.09 ) | 90.27 | 93.77 | 93.13 (+2.86 ) |
| Llama-3.2-3B | 72.55 | 73.84 | 75.06 (+2.51 ) | 45.67 | 51.12 | 53.79 (+8.12 ) | 73.12 | 76.19 | 77.23 (+4.11 ) |
| Llama-3.1-8B | 81.73 | 82.18 | 86.81 (+5.08 ) | 55.62 | 60.79 | 66.17 (+10.55 ) | 77.65 | 85.07 | 86.53 (+8.88 ) |
4 Experiments
4.1 Experimental Setup
Datasets.
Baselines and models.
We conduct of RCR prompting study on ten open-weight models—Qwen (0.5-32B), Llama-3/8B, Mistral-7B, Phi-mini—and two proprietary models, GPT-4o and GPT-4o-mini. We study our DTE framework with 6 models (Qwen 1.5B-14B, Llama-3B and Llama-8B). Baselines are: (i) the single original model; (ii) vanilla MAD with the original MAD prompt Du et al. (2023).
Parameter settings.
During debate we sample each agent once per query at temperature (exploratory) or (deterministic); mixed-teams use one exploratory and two deterministic agents. For evolution we adopt LoRA fine-tuning (rank 128, dropout 0.05) on attention and MLP projections, freezing embeddings and layer norms. GRPO is optimized with AdamW (learning rate , weight decay 0.01, 50-step linear warm-up). Each evolution epoch processes 8k debate traces (2 M tokens) and runs on A100-80 GB GPUs for a 7B model; larger models scale near-linearly.
Evaluation metrics.
Task performance is exact match for GSM-style datasets and accuracy for MC-QA. For RCR evaluation, we also track Sycophancy-Rate: the fraction of agents switching to an incorrect peer answer without adding new reasoning; [incorrect correct] instances during MAD.
4.2 Main Results
Our main results are organized into three main parts: (1) First, we evaluate the effectiveness of Debate–Train–Evolve (DTE) framework, (2) Next, we test its generalization across different reasoning tasks by transferring evolved models to new datasets, and (3) Finally, we analyze the extent of model self-evolution through iterative rounds.
1) Overall DTE performance.
Evolved model using DTE shows an average gain of 8.92% accuracy on GSM-PLUS compared to its vanilla performance. Table 1 contrasts three settings: the single base model (“Original”), a three-agent debate with our RCR prompt (“MAD”), and the evolved single model produced by one Debate–Train–Evolve pass. On GSM-Plus—the hard math dataset—DTE improves every model, with an average gain of +2.38 points over three-agent MAD. Qwen-1.5B shows the largest jump (+13.92 pts), confirming that evolution is most helpful when the base model has head-room and the debate provides diverse traces. On GSM8K the average gain is smaller ( +0.84 pts) because several models were already near their ceiling after debate. ARC-Challenge sees a mixed results: large models benefit (+3.67 pts for Qwen-7B, +8.88 pts for Llama-8B) while small models drift by pt. Overall, DTE shows a mean improvement of 3.06 pts over single model and +1.09 pts over MAD while restoring single-pass inference.
| Model | Fine-tuned on GSM8K | Fine-tuned on GSM-Plus | ||||||
| GSM-Plus | ARC-Easy | ARC-Challenge | CommonsenseQA | GSM8K | ARC-Easy | ARC-Challenge | CommonsenseQA | |
| Qwen-2.5-1.5B | +9.21 | -1.60 | +0.67 | -2.23 | +10.32 | -1.52 | +0.24 | -2.31 |
| Qwen-2.5-3B | +3.79 | +1.27 | +0.83 | +3.26 | +1.36 | +1.09 | +0.60 | +3.26 |
| Qwen-2.5-7B | +1.01 | +1.73 | +4.50 | +3.40 | +1.14 | +1.69 | +3.65 | +3.32 |
| Qwen-2.5-14B | +1.67 | +2.53 | +3.42 | +1.33 | +0.53 | +2.32 | +4.01 | -0.14 |
| Llama-3.2-3B | +6.71 | +2.48 | -1.11 | +3.10 | +3.80 | +1.93 | -3.92 | +3.51 |
| Llama-3.1-8B | +8.13 | +3.91 | +6.74 | +1.10 | +5.15 | +4.88 | +7.84 | +0.85 |
2) Cross-domain generalization.
Our results suggests that DTE improves reasoning that travels beyond the source data, with larger models showing the most stable improvements. Table 2 reports how well the evolved models generalize on other datasets. We test two scenarios: evolve using (i) GSM8K; (ii) GSM-Plus and test on four unseen datasets. When trained on GSM8K, every model gains on GSM-Plus (average +5.8 pts) and on ARC-Challenge (+2.5 pts on average). ARC-Easy also sees small but consistent gains except for the 1.5B model, which drops 1.6 pts. CommonsenseQA improves for 5/6 models, indicating that the reward shaped from mathematical traces still helps improve on commonsense reasoning. Negative deltas are confined to the smallest model (Qwen-1.5B) and to a lesser degree Qwen-3B, suggesting that small models struggles to reconcile new skills with prior knowledge. In contrast, models B never lose more than 0.2 pts on any transfer task. Training on GSM-Plus and testing on GSM8K yields similar behaviour: large gains on the GSM8K (+3.7 pts on average) and moderate gains on others. The symmetry suggests that DTE learns general reasoning heuristics (e.g. numeric decomposition, unit tracking) rather than memorising dataset-specific patterns.
3) How far can a model evolve?
Results show that one evolution round captures nearly all of the available gains. Figure 2 reports accuracy over two evolution rounds for five models on GSM8K and GSM-Plus. Round 1 almost always helps: the smallest model (Qwen-1.5B) jumps from 42.0 55.9 on GSM-Plus and 62.8 73.1 on GSM8K, while Llama-8B gains 10.6 and 5.1 points on the same datasets. The only counter-example is Qwen-7B, which drops 2.4 points on GSM8K despite improving 6.1 on GSM-Plus; upon manual inspection we see that its Round-1 traces over-emphasise shortcut heuristics that hurt easier questions. In Round 2, we observe little improvement and sometimes the performance even drops. Large models ( B) add at most +0.8 points, for Qwen-3B on GSM8K, and more often lose 0.4–1.4 points. The 1.5B model gives back 0.9 points on GSM8K and 2.8 on GSM-Plus, but still ends well above its starting point. Across all runs the mean forgetting is 0.92 pts for models B and 1.6 pts for smaller ones, confirming that smaller models suffers from catastrophic forgetting.
4.3 Ablation Studies
1) Effectiveness of the RCR prompt in MAD.
RCR prompting substantially boost performance over original MAD prompting Du et al. (2023). Figure 3 compares single-model inference, the original debate prompt (MAD@3), and our Reflect–Critique–Refine (RCR-MAD@3) prompt. Across eight diverse models the RCR prompting raises three-agent accuracy by an average of +1.9 pts on GSM8K, +3.7 pts on GSM-Plus, and +0.7 pts on ARC-Challenge. The gain scales with task difficulty: GSM-Plus, which contains harder adversarial questions, benefits the most (up to +7.9 pts for Qwen-1.5B and +6.1 pts for Qwen-7B). On ARC-Challenge improvements are smaller but still positive for 6/8 models. RCR prompting also significantly reduces sycophancy. It halves the mean sycophancy rate (from 0.28 to 0.13 on GSM-Plus) and narrows the verbosity gap by 43 %, indicating that agents now switch answers only when they can articulate a new reasoning step. These observations confirm that RCR is a necessary pre-step for producing high-quality traces later utilized by the DTE training loop.
2) How many agents are enough?
Results shows that three agents MAD captures 85-95 % of the maximum gains. Figure 4 sweeps the agents size from and reports trends on four benchmark. We observe three clear patterns here: 1) Beyond 3-agent the curve plateaus and even oscillates, suggesting the marginal information added by the 4th or 5th agent. 2) Small models benefit most from extra agents. Already strong single-agent (Qwen-14B) adds minimal improvement upon scaling up after three. 3) Harder tasks need (slightly) more agents. On GSM-Plus the optimum often shifts to four or five agents: Qwen-7B reaches its peak accuracy (76.0%) at 7 agents, 1.04 pts above the three-agent setting. ARC-Easy, a much easier dataset, saturates at 2 agents for every model; extra debaters add noise rather than insight.
3) Does agent diversity matter?
We observe two consistent trends here: First, when the individual agents have comparable standalone accuracy, cross-family mixtures beat homogeneous agents team, supporting the idea that architectural diversity yields complementary reasoning paths. Second, when the pool mixes a strong and a weaker model, the debate result gravitates toward the stronger member—adding the weaker agent neither helps nor seriously harms, suggesting that diversity only helps when all agents can contribute novel insights. Complete results for every dataset and roster is available in Appendix B.
4) Why GRPO over other fine-tuning methods?
GRPO consistently outperforms the alternatives, indicating that its relative-advantage reward balances exploration and policy stability better than plain maximum-likelihood (SFT) or preference-only (DPO/PPO) updates. Table 3 compare three update rules under a fixed compute budget: (1) classical supervised fine-tuning on debate answers (SFT); (2) Direct Preference Optimisation using the majority vote as the preferred sample; (3) Group Relative Policy Optimisation (GRPO). GRPO delivers the largest accuracy jump on GSM-Plus for every model size. Both SFT and DPO give smaller gains and even slight regressions on the 3 B model, highlighting the risk of over-fitting when the reward ignores policy shift. We also observe that GRPO keeps across sizes, whereas DPO averages 0.43. The relative-advantage term in GRPO therefore not only boosts reward but also constrains drift, reducing catastrophic forgetting.
| Model | Original (GSM-Plus) | SFT | DPO | GRPO |
| Qwen-2.5-1.5B | 55.92 | |||
| Qwen-2.5-3B | 69.50 | |||
| Qwen-2.5-7B | 74.71 |
5) Data selection strategy.
We test three data sampling schemes on GSM-Plus: Random-2K selects 2000 examples uniformly from the full pool (10552); Debate-Only keeps only data points where agents entered at least one critique round (); All-Traces trains on the entire cleaned set. Table 4 shows that accuracy rises monotonically with coverage: the full corpus beats Debate-Only by 4.43 pts (avg) and Random-2K by 9.17 pts (avg). The gap is largest for Qwen-1.5B, suggesting that smaller models benefit from easier “round-0” examples that Random-2K may miss and Debate-Only discards. We therefore use the full trace set in all other experiments.
| Model | Random-2K | Debate-Only | All-Traces |
| Qwen-1.5B | 44.82 | 51.61 | 55.92 |
| Qwen-3B | 58.10 | 62.70 | 69.50 |
| Qwen-7B | 69.71 | 72.53 | 74.71 |
6) How long do we train?
Figure 5 plots GSM-Plus accuracy as we grow the number of GRPO training steps from 2K to 10K. All models share the similiar trend: rapid gains up to about 8K steps followed by saturation. Small and mid-size models profit the most from the early updates—Qwen-1.5B climbs 8.0 pts between 2K and 6K samples—whereas larger models such as Qwen-14B rise more slowly but steady. Beyond 8K the curve flattens: the average improvement from 8K 10 k is only +0.32 pts while wall-clock time grows by 25%.
7) Does iterative fine-tuning hurt?
Figure 6 plots GSM8K and GSM-Plus accuracy for Qwen-1.5B after the first and second evolution rounds under four sampling temperatures. When we keep the original exploratory setting () the model loses 2.0 pts on GSM8K and gains only 13.5 pts on GSM-Plus—well below the +33.5 pts it achieved in Round 1—confirming a clear case of catastrophic forgetting. Lowering the temperature stabilises training: at Round-2 accuracy is within 0.9 pts of Round 1 on GSM-Plus and almost fully recovers on GSM8K; a deterministic schedule () even adds +3.3 pts on GSM8K but plateaus on GSM-Plus.
The mechanism is visible in the KL divergence between successive students. At we measure for Qwen-1.5B, whereas cuts this to 0.19 and to 0.11, matching the reduction in forgetting. We therefore adopt a linear decay from 0.7 in Round 1 to 0.3 in later rounds for all models up to 3B parameters; larger models did not require temperature adjustment.
5 Conclusion
In this paper, we introduced the Debate, Train, Evolve (DTE) framework, a novel approach enabling language models to autonomously enhance their reasoning capabilities by leveraging multi-agent debate traces. Our Reflect-Critique-Refine prompting strategy significantly improved debate quality, reducing sycophancy and reasoning errors. Experiments demonstrated substantial accuracy gains, notably an average improvement of 8.92% accuracy on the challenging GSM-PLUS dataset. Additionally, we showed strong cross-domain generalization, confirming that our approach captures general reasoning skills rather than dataset-specific patterns. Importantly, DTE effectively combines the benefits of multi-agent debate with the computational efficiency of single-model inference.
Limitations
Despite its effectiveness, our approach has certain limitations. Firstly, iterative fine-tuning within the DTE framework can cause catastrophic forgetting, particularly evident in smaller language models (<3B parameters), leading to potential model collapse. Although we explored several mitigation strategies, completely eliminating this issue remains challenging. Secondly, our framework assumes the availability of high-quality initial debate traces; thus, its efficacy may degrade if debates are of poor quality or if initial agent performance is weak. Third, our study primarily focused on structured reasoning tasks like mathematical and commonsense reasoning. The applicability and effectiveness of DTE on less structured or more open-ended tasks, such as natural language generation or dialogue systems, require further investigation. Lastly, although computationally efficient compared to traditional MAD setups, DTE still incurs higher training costs than standard single-model fine-tuning. Future work should aim to optimize the framework further, enhancing its practicality and accessibility.
Ethics Statement
This study explore the self-evolution of language models using publicly available benchmarks and datasets such as GSM8K, ARC, and CommonsenseQA. All data used in our experiments are non-sensitive and freely accessible, ensuring compliance with ethical research standards and reproducibility. Our method involves fine-tuning on model-generated content, without introducing or relying on any human-annotated private data.
Acknowledgements
This work was supported by NSF NAIRR Pilot with PSC Neocortex, NCSA Delta; Amazon, Cisco Research, Commonwealth Cyber Initiative, Amazon–Virginia Tech Center for Efficient and Robust Machine Learning, and Sanghani Center for AI and Data Analytics at Virginia Tech. The views, findings, conclusions, and recommendations expressed in this work are those of the authors and do not necessarily reflect the opinions of the funding agencies.
Appendix A Datasets Details
We evaluate our Multi-Agent Debate (MAD) approach on five diverse reasoning benchmarks. In the following, we briefly describe each dataset along with their splits. In this paper, we use the test split to evaluate all Small Language Models (SLMs). Table 5 summarizes the splits for each dataset.
| Dataset | Train | Validation | Test |
| GSM8K | 7,473 | – | 1,319 |
| GSM+ | – | 10,552 | 2,400 |
| ARC-Easy | 2,251 | 570 | 2,376 |
| ARC-Challenge | 1,119 | 299 | 1,172 |
| CommonsenseQA | 9,741 | 1,221 | 1,140 |
GSM8K
Cobbe et al. (2021) is a collection of high-quality grade school math word problems that require multi-step reasoning. In the main configuration, the dataset contains a total of 8,790 examples, with 7,473 examples in the training split and 1,319 examples in the test split. These problems typically require between 2 and 8 steps to solve, making it an excellent benchmark for evaluating mathematical reasoning capabilities.
GSM-Plus
Li et al. (2024) extends the GSM8K benchmark with more challenging and diverse mathematical word problems. GSM+ problems generally require more sophisticated multi-step reasoning and often involve more complex mathematical concepts than those in the original GSM8K dataset.
ARC
Clark et al. (2018) comprises two subsets of multiple-choice science questions:
ARC-Easy: Contains 2,251 train, 570 validation, and 2,376 test examples. These questions are answerable by most middle school students and test basic science knowledge.
ARC-Challenge: Contains 1,119 train, 299 validation, and 1,172 test examples. These questions are more challenging and typically answered incorrectly by both retrieval-based algorithms and word co-occurrence algorithms.
CommonsenseQA
Talmor et al. (2019) requires commonsense reasoning to answer multiple-choice questions. It has 9,741 training examples, 1,221 validation examples, and 1,140 test examples. The questions are specifically designed to test commonsense knowledge and reasoning capabilities that go beyond simple factual recall.
Appendix B Implementation Details
Training
For model fine-tuning, we used GRPO to enhance Language Models on the targeted reasoning tasks. Our training pipeline utilized the Unsloth 111https://github.com/unslothai/unsloth and TRL 222https://github.com/huggingface/trl libraries for efficient parameter-efficient fine-tuning with QLoRA Dettmers et al. (2023). Models were trained with a LoRA Hu et al. (2021) rank of 128 and target modules including query, key, value, output, gate, up, and down projections. We used 8-bit Adam optimizer with beta parameters of (0.9, 0.99) and a weight decay of 0.1. The learning rate was set to 5e-6 with a cosine decay schedule and 10% warmup ratio. Training proceeded for 10,000 steps with a per-device batch size of 8. To improve output formatting, we implemented a multi-component reward function consisting of: (1) an answer correctness reward, (2) format adherence rewards for XML tags structure, (3) a numeric response reward, and (4) a tag-counting reward to incentivize proper tag usage. Each model was instructed to output responses in a structured XML format with separate <reasoning> and <answer> tags to facilitate consistent answer extraction and evaluation. To manage memory constraints on high-end GPUs, we set the maximum sequence length to 2048 tokens with a 512-token maximum for prompts and 1536 tokens for model completions.
Inference
We conducted all model inferences using NVIDIA H100-80GB, A100-80GB, L40-48GB, and A40-48GB GPUs. For efficient inference, we used the vLLM library Kwon et al. (2023) 333https://docs.vllm.ai/en/latest/, dynamically allocating the required number of GPUs to load each model. Multi-GPU utilization was enabled using Hugging Face Accelerate 444https://github.com/huggingface/accelerate for model sharding and speed optimization.
Appendix C Reflect–Critique–Refine Prompt Design
Appendix D Additional Self-Evolution Results
In this section, we present a comprehensive analysis of our Debate, Train, Evolve framework across multiple experimental settings. We first examine the impact of various GRPO configurations, followed by analyses of multi-round training effects, and finally cross-domain generalization results. Our experiments utilize a diverse set of models ranging from 1.5B to 14B parameters and evaluate performance on challenging reasoning benchmarks including GSM8K, GSM-Plus, ARC-Challenge, ARC-Easy, and CommonsenseQA.
D.1 Complete GRPO results (all steps, temperature)
We begin by investigating how different GRPO hyperparameters affect model performance. Tables 6, 7, and 8 present results across three datasets (GSM8K, GSM-Plus, and ARC-Challenge) for six different model configurations, varying training steps (2000, 5000, and 10000) and sampling temperatures (0.8 and 0.2).
Several key patterns emerge from these results. First, we observe that larger models (7B+) generally maintain or improve their performance through GRPO fine-tuning, while smaller models (particularly Llama-3B) occasionally exhibit catastrophic forgetting at higher step counts. Second, lower temperature (0.2) typically yields more stable optimization trajectories for most model configurations, especially at higher step counts. This supports our hypothesis that constraining policy drift during fine-tuning is crucial for successful reasoning evolution.
Notably, the Qwen-2.5-3B model demonstrates remarkable stability across configurations, with consistent performance gains on GSM-Plus (from 61.75% to 69.50%) and robust maintenance of GSM8K performance. In contrast, the Llama-3B model shows significant performance degradation at higher step counts with 0.8 temperature, dropping to near-random performance (2.73%) after 10000 steps on GSM8K, while maintaining better stability at 0.2 temperature.
For ARC-Challenge, we observe that all models benefit from MAD evolution, with particularly strong gains for Qwen-2.5-7B (from 87.22% to 91.64%) and Llama-8B (from 77.65% to 85.07%). These results suggest that our framework effectively generalizes across both mathematical reasoning and scientific question-answering domains.
D.2 Complete Round 2 MAD Results
After the first round of GRPO fine-tuning, we evaluated the performance of models in a multi-agent debate setting to assess how evolution affects collaborative reasoning. Table 10 presents these results across different debate configurations: exponential temperature scaling (Exp), default settings (Default), temperature-4 settings (temp4), and deterministic setting (Det).
The MAD Round 2 results demonstrate that evolved models generally maintain their collaborative reasoning capabilities after GRPO fine-tuning. For most models, MAD performance after evolution either improves or remains comparable to the original MAD results. The Qwen-2.5-7B model, for instance, achieves 77.75% accuracy on GSM-Plus under the temp4 configuration, which represents a 3.58% improvement over its original MAD performance.
Interestingly, we observe that different debate configurations yield varying results across model sizes. Smaller models like Qwen-2.5-1.5B show significant performance variation across configurations, with deterministic settings yielding the best results (69.07% on GSM8K and 56.62% on GSM-Plus). In contrast, larger models like Qwen-2.5-7B demonstrate more consistent performance across configurations.
The exponential temperature scaling configuration generally underperforms other settings, particularly for smaller models. This suggests that controlled diversity in debate is beneficial, but excessive exploration may hinder collaborative reasoning effectiveness.
D.3 GRPO round 2 results
To investigate the effects of iterative evolution, we conducted a second round of GRPO fine-tuning on models that had already undergone one round of evolution. Table 9 presents these results for four model configurations across two datasets (GSM8K and GSM-Plus).
The second round of GRPO training reveals interesting dynamics in model evolution. For the Qwen family of models, we observe continued performance improvements or stability across most configurations. The Qwen-2.5-7B model, for instance, achieves further gains on GSM-Plus, reaching 73.75% accuracy (a 5.13% improvement over its first round GRPO performance).
However, the Llama-3B model exhibits significant performance degradation in certain configurations, particularly at higher step counts with 0.8 temperature (dropping to 35.63% on GSM8K and 23.02% on GSM-Plus). This reinforces our finding that smaller models are more sensitive to optimization instability during iterative fine-tuning. Importantly, using a lower temperature of 0.2 substantially mitigates this issue, allowing the Llama-3B model to maintain competitive performance (73.62% on GSM8K) even after two rounds of evolution.
These results highlight the importance of careful hyperparameter selection during iterative self-evolution, particularly for smaller models that may be more susceptible to catastrophic forgetting or excessive policy drift.
D.4 Complete Round 3 MAD Results
To investigate the long-term stability of collaborative reasoning capabilities through multiple evolution iterations, we conducted a third round of multi-agent debate after the second round of GRPO fine-tuning. Table 11 presents these results for three Qwen models across the same four debate configurations.
The Round 3 MAD results reveal interesting trends in iterative evolution. For the Qwen-2.5-3B and Qwen-2.5-7B models, performance remains relatively stable across debate configurations, indicating robust retention of reasoning capabilities through multiple fine-tuning iterations. However, the Qwen-2.5-1.5B model shows more variable performance, particularly under the exponential temperature scaling configuration where it drops to 44.28% on GSM8K.
Notably, the deterministic debate setting (Det) consistently produces the best or near-best performance across all models and datasets, suggesting that reduced randomness in collaborative reasoning becomes increasingly important after multiple evolution rounds. This aligns with our hypothesis that controlling policy drift is crucial for successful iterative evolution.
The stability of larger models (3B+) across multiple evolution rounds indicates that our Debate, Train, Evolve framework can support continuous improvement without substantial performance degradation when applied to sufficiently capable base models.
D.5 Complete Cross Domain Task Results
A key question for self-evolution frameworks is whether improvements generalize beyond the training domain. Table 12 presents results for models fine-tuned on either GSM8K or GSM-Plus and evaluated on multiple out-of-domain tasks including ARC-Easy, ARC-Challenge, and CommonsenseQA.
The cross-domain results reveal impressive generalization capabilities. Models fine-tuned on mathematical reasoning tasks (GSM8K and GSM-Plus) show substantial performance improvements not only on the alternative math dataset but also on science and commonsense reasoning benchmarks. For instance, the Qwen-2.5-14B model fine-tuned on GSM8K achieves 98.19% accuracy on ARC-Easy, 93.69% on ARC-Challenge, and 83.70% on CommonsenseQA.
Interestingly, models fine-tuned on GSM-Plus generally perform better on GSM8K than vice versa. For example, the Qwen-2.5-1.5B model achieves 73.09% on GSM8K when fine-tuned on GSM-Plus, but only 51.21% on GSM-Plus when fine-tuned on GSM8K. This asymmetry suggests that GSM-Plus may require more diverse reasoning strategies that transfer well to simpler tasks.
The strong cross-domain performance demonstrates that our Debate, Train, Evolve framework does not simply optimize for specific datasets but instead enhances fundamental reasoning capabilities that generalize across tasks. This is a critical advantage over traditional supervised fine-tuning approaches that often exhibit limited transferability.
| Model | Base Performance | MAD | GRPO (Temperature 0.8) | GRPO (Temperature 0.2) | |||||
| Train | Test | 2k steps | 5k steps | 10k steps | 2k steps | 5k steps | 10k steps | ||
| Qwen-2.5-1.5B | 81.55 | 62.77 | 72.33 | 67.78 | 71.42 | 71.04 | 73.09 | 66.49 | 53.98 |
| Qwen-2.5-3B | 91.28 | 84.08 | 85.14 | 85.06 | 85.14 | 86.13 | 84.00 | 86.05 | 84.38 |
| Qwen-2.5-7B | 94.29 | 90.67 | 91.21 | 88.32 | 86.73 | 84.00 | 86.96 | 86.35 | 88.02 |
| Llama-3B | 83.90 | 72.55 | 73.84 | 69.22 | 21.53 | 2.73 | 72.40 | 75.06 | 3.26 |
| Llama-8B | 89.08 | 81.73 | 82.18 | 84.61 | 85.29 | 85.22 | 86.81 | 84.91 | 0.15 |
| Qwen-2.5-14B | 94.89 | 92.80 | 93.33 | 87.72 | 89.84 | 91.81 | 86.58 | 89.34 | 93.74 |
| Model | Base Performance | MAD | GRPO (Temperature 0.8) | GRPO (Temperature 0.2) | |||||
| Train | Test | 2k steps | 5k steps | 10k steps | 2k steps | 5k steps | 10k steps | ||
| Qwen-2.5-1.5B | 42.40 | 42.00 | 51.62 | 47.49 | 54.46 | 19.00 | 52.33 | 53.04 | 55.92 |
| Qwen-2.5-3B | 61.14 | 61.75 | 67.79 | 66.21 | 66.71 | 69.13 | 64.04 | 67.25 | 68.25 |
| Qwen-2.5-7B | 68.27 | 68.62 | 74.17 | 64.71 | 73.38 | 74.71 | 67.75 | 72.54 | 74.50 |
| Llama-3B | 47.68 | 45.67 | 51.12 | 52.38 | 53.29 | 52.33 | 51.79 | 49.54 | 53.79 |
| Llama-8B | 58.56 | 55.62 | 60.79 | 64.96 | 61.58 | 66.17 | 65.08 | 63.46 | 60.46 |
| Qwen-2.5-14B | 71.11 | 71.79 | 77.25 | 70.79 | 73.54 | 75.88 | 73.00 | 73.42 | 75.62 |
| Model | Base Performance | MAD | GRPO (Temperature 0.8) | GRPO (Temperature 0.2) | |||||
| Train | Test | 2k steps | 5k steps | 10k steps | 2k steps | 5k steps | 10k steps | ||
| Qwen-2.5-1.5B | — | 69.21 | 68.52 | 30.03 | 62.63 | 68.36 | 47.27 | 51.88 | 67.51 |
| Qwen-2.5-3B | — | 83.53 | 84.64 | 81.66 | 80.29 | 83.63 | 81.91 | 79.78 | 83.95 |
| Qwen-2.5-7B | — | 87.22 | 91.64 | 88.57 | 88.48 | 90.63 | 88.43 | 88.57 | 90.89 |
| Llama-3B | — | 73.12 | 76.19 | 75.51 | 74.32 | 76.87 | 76.79 | 74.57 | 77.23 |
| Llama-8B | — | 77.65 | 85.07 | 83.70 | 84.45 | 86.03 | 84.98 | 85.53 | 86.53 |
| Qwen-2.5-14B | — | 90.27 | 93.77 | 91.81 | 92.49 | 93.13 | 91.47 | 91.47 | 92.67 |
| Model | Dataset | GRPO Round 2 (Temp 0.8) | GRPO Round 2 (Temp 0.2) | ||
| 2k steps | 5k steps | 2k steps | 5k steps | ||
| Qwen-2.5-1.5B | GSM8K | 65.73 | 68.54 | 69.98 | 72.18 |
| GSM-Plus | 47.38 | 50.12 | 46.37 | 48.04 | |
| Qwen-2.5-3B | GSM8K | 84.84 | 86.05 | 84.46 | 84.08 |
| GSM-Plus | 65.71 | 67.96 | 65.67 | 67.00 | |
| Qwen-2.5-7B | GSM8K | 86.28 | 87.19 | 88.17 | 87.34 |
| GSM-Plus | 69.42 | 73.75 | 70.54 | 73.12 | |
| Llama-3B | GSM8K | 55.88 | 35.63 | 73.62 | 64.29 |
| GSM-Plus | 48.75 | 23.02 | 52.42 | 25.08 | |
| Model | Dataset | MAD Configuration | |||
| Exp | Default | temp4 | Det | ||
| Qwen-2.5-1.5B | GSM8K | 46.32 | 66.34 | 68.61 | 69.07 |
| GSM-Plus | 22.09 | 53.18 | 55.62 | 56.62 | |
| Qwen-2.5-3B | GSM8K | 84.08 | 86.66 | 86.35 | 86.50 |
| GSM-Plus | 69.62 | 70.25 | 69.67 | 70.29 | |
| Qwen-2.5-7B | GSM8K | 91.36 | 90.75 | 91.05 | 89.99 |
| GSM-Plus | 76.42 | 77.00 | 77.75 | 77.62 | |
| Llama-3B | GSM8K | 66.26 | 75.97 | 75.51 | 75.36 |
| GSM-Plus | 53.62 | 54.58 | 55.96 | 56.04 | |
| Llama-8B | GSM8K | 84.69 | 85.90 | 86.96 | 85.60 |
| GSM-Plus | 65.00 | 65.92 | 66.46 | 66.50 | |
| Model | Dataset | MAD Configuration | |||
| Exp | Default | temp4 | Det | ||
| Qwen-2.5-1.5B | GSM8K | 44.28 | 60.65 | 67.70 | 72.40 |
| GSM-Plus | 35.54 | 48.62 | 51.67 | 51.75 | |
| Qwen-2.5-3B | GSM8K | 83.78 | 85.60 | 85.75 | 86.13 |
| GSM-Plus | 63.67 | 63.42 | 64.16 | 64.47 | |
| Qwen-2.5-7B | GSM8K | 89.76 | 91.05 | 90.90 | 91.13 |
| GSM-Plus | 69.67 | 69.85 | 70.50 | 69.88 | |
| Model | Fine-tuned on | Evaluation Dataset | ||||
| GSM8K | GSM-Plus | ARC-Easy | ARC-Challenge | CommonsenseQA | ||
| Qwen-2.5-1.5B | GSM8K | — | 51.21 | 85.02 | 69.88 | 64.29 |
| GSM-Plus | 73.09 | — | 85.10 | 69.45 | 64.21 | |
| Qwen-2.5-3B | GSM8K | — | 65.54 | 93.94 | 84.30 | 75.92 |
| GSM-Plus | 86.50 | — | 94.15 | 84.13 | 75.92 | |
| Qwen-2.5-7B | GSM8K | — | 69.63 | 96.42 | 91.72 | 82.96 |
| GSM-Plus | 91.81 | — | 96.38 | 90.87 | 82.88 | |
| Llama-3B | GSM8K | — | 52.38 | 87.12 | 72.01 | 68.14 |
| GSM-Plus | 76.35 | — | 86.57 | 69.20 | 68.55 | |
| Llama-8B | GSM8K | — | 63.75 | 93.01 | 84.39 | 74.12 |
| GSM-Plus | 86.88 | — | 93.98 | 85.49 | 73.87 | |
| Qwen-2.5-14B | GSM8K | — | 73.46 | 98.19 | 93.69 | 83.70 |
| GSM-Plus | 93.33 | — | 97.98 | 94.28 | 82.23 | |
Appendix E Complete Results of Large-scale Empirical Study on MAD using RCR Prompting
This section presents a comprehensive analysis of our large-scale empirical investigation into Multi-Agent Debate (MAD) using Recursive Critical Reflection (RCR) prompting across five diverse benchmarks: GSM8K, GSM-Plus, ARC-Easy, ARC-Challenge, and CommonsenseQA. Through extensive experimentation involving various model combinations and parameter settings, we evaluate how collaborative reasoning among multiple language model agents affects problem-solving performance.
E.1 Evaluation Metrics and Methodology
To facilitate systematic comparison and analysis of debate outcomes, we track the following key metrics across all debate configurations:
Accuracy: The primary performance measure, representing the percentage of problems correctly solved after the debate process concludes.
(Performance Delta): Measures the performance change relative to appropriate baselines. We report several variants including:
(vs Base): Change compared to the single agent’s performance
(vs Lower Agent): Change compared to the lower-performing agent in cross-agent debates
(vs Upper Agent): Change compared to the better-performing agent in cross-agent debates
(vs Lowest): Change compared to the lowest-performing agent in three-agent settings
Debate Rounds: The average number of interaction rounds required to reach consensus or the maximum allowed limit, indicating debate efficiency.
Sycophancy: A normalized measure (per data points) quantifying the tendency of agents to abandon their answers in favor of matching another agent’s previous response, providing insights into social influence dynamics.
State Transitions: Tracked as CI (correct to incorrect) and IC (incorrect to correct) counts, these reveal the qualitative nature of answer changes during debate.
Debate Helped: The overall count of instances where the debate process improved the final outcome compared to initial responses.
Our evaluation spans multiple dimensions of agent configuration:
Agent Settings: We systematically vary temperature parameter across four settings:
Default: Balanced temperature
Deterministic (Det.): Lower temperature for more consistent outputs
Exploratory (Exp.): Higher temperature for more diverse responses
Mixed: Combinations of the above settings across different agents
Debate Structures: We investigate four primary debate configurations:
Single-Model Debate: Multiple instances of the same model with varied parameter settings
Cross-Agent Debate: Two different models debating with various parameter settings
Three Identical Agents: Three instances of the same model with potentially different settings
Three Varied Agents: Three different models engaging in debate
E.2 Overview of Results Organization
Our extensive experimental results are organized in Tables 13-32, systematically covering all five datasets with the four debate configurations described above. For each dataset, we present:
Table set 1 (Tables 13-16): Performance on GSM8K
Table set 2 (Tables 17-20): Performance on GSM-Plus
Table set 3 (Tables 21-24): Performance on ARC-Easy
Table set 4 (Tables 25-28): Performance on ARC-Challenge
Table set 5 (Tables 29-32): Performance on CommonsenseQA
E.3 Key Findings and Patterns
E.3.1 Impact of Agent Settings
Our analysis reveals that agent parameter settings significantly influence debate outcomes across all datasets. We observe that while the Default setting provides reliable performance, Exploratory settings often lead to higher variance in outcomes, sometimes yielding exceptional improvements but also risking performance degradation. The Deterministic setting generally produces more consistent but potentially conservative results.
The sycophancy metric proves particularly informative, showing higher values in debates between models with substantial performance gaps. This suggests that lower-performing models tend to defer to higher-performing ones, which can be either beneficial or detrimental depending on the initial state distribution.
E.3.2 Cross-Model Debate Dynamics
In cross-agent debates (Tables 10-14), we find that pairing models with complementary strengths often produces synergistic effects. The metrics relative to both upper and lower agents reveal important patterns: when a high-performing model debates with a weaker one, the debate outcome typically falls between their individual performances but closer to the stronger model’s baseline.
State transitions (CI and IC) provide valuable insights into debate quality. A high IC rate coupled with a low CI rate indicates constructive debate where correct reasoning prevails, while the opposite pattern signals problematic dynamics where convincing but incorrect reasoning dominates.
E.3.3 Three-Agent Debate Effectiveness
The introduction of a third agent creates more complex interaction patterns. Three-agent debates consistently show lower sycophancy rates compared to two-agent settings, suggesting that the presence of multiple perspectives reduces blind conformity. When all three agents are identical, we observe that diversity in parameter settings typically outperforms homogeneous settings.
In three varied agent debates, we find particularly interesting results when combining models of different sizes and architectures. As shown in Table 16, certain combinations like "Qwen-2.5-3B + Phi-mini-3.8B + Llama-3.1-3B" achieve accuracy improvements even compared to the highest-performing individual agent, suggesting effective complementarity between these models’ reasoning approaches.
E.3.4 Dataset-Specific Patterns
Our results indicate substantial variation in debate effectiveness across different datasets:
GSM8K and GSM+: Harder Mathematical reasoning tasks (GSM-Plus) show the most consistent benefits from debate, with average debate rounds typically higher than other datasets, suggesting that step-by-step verification is particularly valuable for these problems.
ARC-Easy and ARC-Challenge: Multiple-choice science questions reveal interesting patterns where sycophancy is generally lower, but debate can still improve performance when appropriately configured.
CommonsenseQA: This dataset exhibits unique characteristics where debates tend to conclude more quickly, suggesting that commonsense reasoning may be less amenable to explicit verification through debate.
E.4 Conclusion
Tables 13-32 collectively present a comprehensive empirical foundation for understanding the effects of Multi-Agent Debate using RCR prompting across diverse reasoning tasks. The metrics reveal nuanced patterns in how debate influences performance, with clear evidence that appropriate configuration of debate participants and settings can yield substantial improvements over single-agent performance.
The consistent tracking of accuracy, deltas, debate rounds, sycophancy, and state transitions provides a multi-dimensional view of debate quality beyond simple performance measures. These results demonstrate that MAD is not universally beneficial but rather depends critically on the specific combination of models, parameter settings, and problem domains. Our findings establish an important baseline for future research on collaborative reasoning between language models, highlighting both the potential and the challenges of multi-agent approaches to complex problem-solving.
Agent 1 Agent 2 Agent Settings MAD Accuracy Debate Sycophancy CI IC Debate (RCR Prompting) Rounds (Avg / 1319) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Default 47.38 5.38 251 220 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Deterministic 47.31 5.31 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Exploratory 39.20 2.8 274 234 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Det. & Exp. 43.14 1.14 262 226 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Default 70.89 8.12 352 317 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Deterministic 63.46 0.69 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Exploratory 71.57 8.8 449 399 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Det. & Exp. 72.33 9.56 423 377 Qwen-2.5-3B Qwen-2.5-3B Both: Default 86.05 0.91 115 104 Qwen-2.5-3B Qwen-2.5-3B Both: Deterministic 84.99 0.15 0 0 Qwen-2.5-3B Qwen-2.5-3B Both: Exploratory 85.52 0.38 116 103 Qwen-2.5-3B Qwen-2.5-3B Both: Det. & Exp. 86.28 1.14 106 101 Qwen-2.5-7B Qwen-2.5-7B Both: Default 91.74 1.07 53 49 Qwen-2.5-7B Qwen-2.5-7B Both: Deterministic 90.60 0.07 0 0 Qwen-2.5-7B Qwen-2.5-7B Both: Exploratory 91.21 0.54 59 57 Qwen-2.5-7B Qwen-2.5-7B Both: Det. & Exp. 91.51 0.84 57 55 Qwen-2.5-14B Qwen-2.5-14B Both: Default 93.48 0.68 46 43 Qwen-2.5-14B Qwen-2.5-14B Both: Deterministic 93.18 0.38 0 0 Qwen-2.5-14B Qwen-2.5-14B Both: Exploratory 93.33 0.53 48 48 Qwen-2.5-14B Qwen-2.5-14B Both: Det. & Exp. 93.63 0.83 44 39 Qwen-2.5-32B Qwen-2.5-32B Both: Default 95.00 0.08 21 20 Qwen-2.5-32B Qwen-2.5-32B Both: Deterministic 94.77 0.15 0 0 Qwen-2.5-32B Qwen-2.5-32B Both: Exploratory 95.38 0.46 32 31 Qwen-2.5-32B Qwen-2.5-32B Both: Det. & Exp. 95.30 23 21 Llama-3.1-3B Llama-3.1-3B Both: Default 74.91 2.36 208 183 Llama-3.1-3B Llama-3.1-3B Both: Deterministic 74.37 1.82 0 0 Llama-3.1-3B Llama-3.1-3B Both: Exploratory 72.40 0.15 225 202 Llama-3.1-3B Llama-3.1-3B Both: Det. & Exp. 73.84 1.29 193 175 Llama-3.1-8B Llama-3.1-8B Both: Default 82.56 0.83 116 105 Llama-3.1-8B Llama-3.1-8B Both: Deterministic 81.50 0.23 0 0 Llama-3.1-8B Llama-3.1-8B Both: Exploratory 80.67 1.06 162 149 Llama-3.1-8B Llama-3.1-8B Both: Det. & Exp. 82.18 0.45 142 126 Phi-mini-3.8B Phi-mini-3.8B Both: Default 87.72 0.84 101 95 Phi-mini-3.8B Phi-mini-3.8B Both: Deterministic 86.73 0.15 2 1 Phi-mini-3.8B Phi-mini-3.8B Both: Exploratory 87.95 1.07 112 99 Phi-mini-3.8B Phi-mini-3.8B Both: Det. & Exp. 87.34 0.46 103 95 Mistral-7B Mistral-7B Both: Default 33.74 12.36 454 340 Mistral-7B Mistral-7B Both: Deterministic 20.02 0 0 Mistral-7B Mistral-7B Both: Exploratory 35.71 14.33 509 381 Mistral-7B Mistral-7B Both: Det. & Exp. 33.51 433 334
Agent 1 Agent 2 Agent Settings Accuracy (Lower Agent) (Upper Agent) Debate Sycophancy CI IC Debate Rounds (Avg / 1319) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B 1: Default & 2: Default 62.40 20.4 0.37 434 387 Qwen-2.5-0.5B Qwen-2.5-1.5B 1: Det. & 2: Det. 62.32 20.32 0.45 357 323 Qwen-2.5-0.5B Qwen-2.5-1.5B 1: Exp. & 2: Exp. 58.91 16.91 3.86 531 448 Qwen-2.5-0.5B Qwen-2.5-1.5B 1: Det. & 2: Exp. 60.88 18.88 1.89 416 344 Qwen-2.5-0.5B Qwen-2.5-1.5B 1: Exp. & 2: Det. 61.18 19.18 1.59 474 425 Qwen-2.5-1.5B Llama-3.1-3B 1: Default & 2: Default 76.42 13.65 3.87 388 342 Qwen-2.5-1.5B Llama-3.1-3B 1: Det. & 2: Det. 75.59 12.82 3.04 285 258 Qwen-2.5-1.5B Llama-3.1-3B 1: Exp. & 2: Exp. 76.57 13.8 4.02 416 355 Qwen-2.5-1.5B Llama-3.1-3B 1: Det. & 2: Exp. 75.06 12.29 2.51 362 326 Qwen-2.5-1.5B Llama-3.1-3B 1: Exp. & 2: Det. 76.04 13.27 3.49 383 331 Qwen-2.5-3B Phi-mini-3.8B 1: Default & 2: Default 87.41 2.27 0.53 128 114 Qwen-2.5-3B Phi-mini-3.8B 1: Det. & 2: Det. 85.97 0.83 0.91 82 72 Qwen-2.5-3B Phi-mini-3.8B 1: Exp. & 2: Exp. 88.63 3.49 1.75 155 142 Qwen-2.5-3B Phi-mini-3.8B 1: Det. & 2: Exp. 86.73 1.59 0.15 105 99 Qwen-2.5-3B Phi-mini-3.8B 1: Exp. & 2: Det. 88.10 2.96 1.22 135 126 Qwen-2.5-1.5B Qwen-2.5-3B 1: Default & 2: Default 82.71 19.94 2.43 370 359 Qwen-2.5-1.5B Qwen-2.5-3B 1: Det. & 2: Det. 81.27 18.5 3.87 284 275 Qwen-2.5-1.5B Qwen-2.5-3B 1: Exp. & 2: Exp. 83.17 20.4 1.97 414 392 Qwen-2.5-1.5B Qwen-2.5-3B 1: Det. & 2: Exp. 82.87 20.1 2.27 328 310 Qwen-2.5-1.5B Qwen-2.5-3B 1: Exp. & 2: Det. 82.26 19.49 2.88 384 372 Llama-3.1-3B Llama-3.1-8B 1: Default & 2: Default 78.54 5.99 3.19 213 195 Llama-3.1-3B Llama-3.1-8B 1: Det. & 2: Det. 79.23 6.68 2.5 159 143 Llama-3.1-3B Llama-3.1-8B 1: Exp. & 2: Exp. 77.10 4.55 4.63 238 224 Llama-3.1-3B Llama-3.1-8B 1: Det. & 2: Exp. 79.83 7.28 1.9 211 183 Llama-3.1-3B Llama-3.1-8B 1: Exp. & 2: Det. 77.18 4.63 4.55 183 173 Qwen-2.5-7B Qwen-2.5-14B 1: Default & 2: Default 92.19 1.52 0.61 63 61 Qwen-2.5-7B Qwen-2.5-14B 1: Det. & 2: Det. 92.04 1.37 0.76 53 50 Qwen-2.5-7B Qwen-2.5-14B 1: Exp. & 2: Exp. 93.10 2.43 0.3 72 68 Qwen-2.5-7B Qwen-2.5-14B 1: Det. & 2: Exp. 92.19 1.52 0.61 58 58 Qwen-2.5-7B Qwen-2.5-14B 1: Exp. & 2: Det. 92.80 2.13 64 60
Agent 1 Agent 2 Agent 3 Agent Settings Accuracy (Improvement) Debate Sycophancy CI IC Debate Rounds (Avg / 1319) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B All: Default 0.3 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B All: Deterministic 5.31 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B All: Exploratory 5.91 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 1 Det, 2 Exp 3.64 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 2 Det, 1 Exp 1.06 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B All: Default 9.71 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B All: Deterministic 1.22 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B All: Exploratory 12.36 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 1 Det, 2 Exp 12.06 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 2 Det, 1 Exp 9.48 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B All: Default 1.82 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B All: Deterministic 0.15 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B All: Exploratory 2.5 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 1 Det, 2 Exp 1.59 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 2 Det, 1 Exp 0.91 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B All: Default 2.36 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B All: Deterministic 0.07 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B All: Exploratory 1.75 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 1 Det, 2 Exp 1.45 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 2 Det, 1 Exp 1.29 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B All: Default Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B All: Deterministic Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B All: Exploratory Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 1 Det, 2 Exp Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 2 Det, 1 Exp Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B All: Default Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B All: Deterministic Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B All: Exploratory Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 1 Det, 2 Exp Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 2 Det, 1 Exp Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B All: Default 1.52 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B All: Deterministic 0.22 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B All: Exploratory 1.22 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 1 Det, 2 Exp 0.99 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 2 Det, 1 Exp 0.84 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B All: Default 0.08 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B All: Deterministic 0.61 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B All: Exploratory 0.23 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 1 Det, 2 Exp 1.14 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 2 Det, 1 Exp 0.38 Mistral-7B Mistral-7B Mistral-7B All: Default 16.45 Mistral-7B Mistral-7B Mistral-7B All: Deterministic 1.36 Mistral-7B Mistral-7B Mistral-7B All: Exploratory 17.89 Mistral-7B Mistral-7B Mistral-7B 1 Det, 2 Exp 17.51 Mistral-7B Mistral-7B Mistral-7B 2 Det, 1 Exp 13.95 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B All: Default 2.5 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B All: Deterministic 0.23 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B All: Exploratory 1.97 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 1 Det, 2 Exp 1.59 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 2 Det, 1 Exp 0.53
Agent 1 Agent 2 Agent 3 Agent Settings Accuracy (vs Lowest) Debate Sycophancy CI IC Debate Rounds (Avg / 1319) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-3B All: Default 4.32 Qwen-2.5-0.5B Qwen-2.5-1.5B Llama-3.1-3B All: Default 3.03 Qwen-2.5-0.5B Qwen-2.5-1.5B Phi-mini-3.8B All: Default 10.84 Qwen-2.5-0.5B Qwen-2.5-3B Llama-3.1-3B All: Default 5.99 Qwen-2.5-0.5B Qwen-2.5-3B Phi-mini-3.8B All: Default 3.24 Qwen-2.5-0.5B Llama-3.1-3B Phi-mini-3.8B All: Default 10.31 Qwen-2.5-1.5B Qwen-2.5-3B Llama-3.1-3B All: Default 2.43 Qwen-2.5-1.5B Qwen-2.5-3B Phi-mini-3.8B All: Default 1.66 Qwen-2.5-1.5B Llama-3.1-3B Phi-mini-3.8B All: Default 5.68 Qwen-2.5-3B Phi-mini-3.8B Llama-3.1-3B All: Default 0.08 Qwen-2.5-3B Qwen-2.5-3B Phi-mini-3.8B All: Default 0.76 Qwen-2.5-3B Phi-mini-3.8B Phi-mini-3.8B All: Default 0.91 Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-1.5B All: Default 5.69 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-1.5B All: Default 7.65
Agent 1 Agent 2 Agent Settings MAD Accuracy Debate Sycophancy CI IC Debate (RCR Prompting) Rounds (Avg / 2400) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Default 27.33 2.54 348 295 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Deterministic 29.25 4.46 2 1 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Exploratory 23.12 1.67 351 289 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Det. & Exp. 27.33 2.54 396 336 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Default 53.12 11.12 555 502 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Deterministic 47.29 5.29 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Exploratory 51.62 9.62 647 551 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Det. & Exp. 52.29 10.29 528 477 Qwen-2.5-3B Qwen-2.5-3B Both: Default 67.42 5.67 225 213 Qwen-2.5-3B Qwen-2.5-3B Both: Deterministic 67.38 5.63 0 0 Qwen-2.5-3B Qwen-2.5-3B Both: Exploratory 67.79 6.04 296 265 Qwen-2.5-3B Qwen-2.5-3B Both: Det. & Exp. 66.46 4.71 223 208 Qwen-2.5-7B Qwen-2.5-7B Both: Default 74.17 5.55 135 127 Qwen-2.5-7B Qwen-2.5-7B Both: Deterministic 73.62 5.00 0 0 Qwen-2.5-7B Qwen-2.5-7B Both: Exploratory 74.17 5.55 158 150 Qwen-2.5-7B Qwen-2.5-7B Both: Det. & Exp. 74.46 5.84 126 118 Qwen-2.5-14B Qwen-2.5-14B Both: Default 77.21 5.42 102 100 Qwen-2.5-14B Qwen-2.5-14B Both: Deterministic 76.25 4.46 0 0 Qwen-2.5-14B Qwen-2.5-14B Both: Exploratory 77.25 5.46 128 123 Qwen-2.5-14B Qwen-2.5-14B Both: Det. & Exp. 76.96 5.17 99 93 Qwen-2.5-32B Qwen-2.5-32B Both: Default 73.33 0.87 62 59 Qwen-2.5-32B Qwen-2.5-32B Both: Deterministic 72.79 0.33 0 0 Qwen-2.5-32B Qwen-2.5-32B Both: Exploratory 73.42 0.96 91 88 Qwen-2.5-32B Qwen-2.5-32B Both: Det. & Exp. 73.46 1.00 70 68 Phi-mini-3.8B Phi-mini-3.8B Both: Default 69.62 6.20 204 191 Phi-mini-3.8B Phi-mini-3.8B Both: Deterministic 69.21 5.79 6 3 Phi-mini-3.8B Phi-mini-3.8B Both: Exploratory 70.38 6.96 267 242 Phi-mini-3.8B Phi-mini-3.8B Both: Det. & Exp. 69.42 6.00 203 188 Mistral-7B Mistral-7B Both: Default 23.42 8.38 576 434 Mistral-7B Mistral-7B Both: Deterministic 14.33 0.71 4 2 Mistral-7B Mistral-7B Both: Exploratory 23.29 8.25 586 437 Mistral-7B Mistral-7B Both: Det. & Exp. 22.75 7.71 556 414 Llama-3.1-3B Llama-3.1-3B Both: Default 51.58 5.91 439 378 Llama-3.1-3B Llama-3.1-3B Both: Deterministic 50.50 4.83 0 0 Llama-3.1-3B Llama-3.1-3B Both: Exploratory 51.12 5.45 482 406 Llama-3.1-3B Llama-3.1-3B Both: Det. & Exp. 50.75 5.08 381 333 Llama-3.1-8B Llama-3.1-8B Both: Default 62.04 6.42 313 274 Llama-3.1-8B Llama-3.1-8B Both: Deterministic 61.04 5.42 0 0 Llama-3.1-8B Llama-3.1-8B Both: Exploratory 60.79 5.17 340 303 Llama-3.1-8B Llama-3.1-8B Both: Det. & Exp. 60.96 5.34 304 273
Agent 1 Agent 2 Agent Settings MAD Lower Upper Debate Sycophancy CI IC Debate Accuracy Rounds (Avg / 2400) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Default 16.59 0.62 314 628 548 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Deterministic 17.88 0.67 292 565 505 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Exploratory 14.75 2.46 320 722 604 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Det. & Exp. 15.25 1.96 301 588 492 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Exp. & Det. 19.46 2.25 278 750 664 Qwen-2.5-1.5B Llama-3.1-3B Both: Default 12.42 8.75 232 612 532 Qwen-2.5-1.5B Llama-3.1-3B Both: Deterministic 12.37 8.70 224 489 435 Qwen-2.5-1.5B Llama-3.1-3B Both: Exploratory 12.21 8.54 255 696 602 Qwen-2.5-1.5B Llama-3.1-3B Both: Det. & Exp. 11.29 7.62 249 555 488 Qwen-2.5-1.5B Llama-3.1-3B Both: Exp. & Det. 12.58 8.91 249 603 533 Qwen-2.5-3B Phi-mini-3.8B Both: Default 8.46 6.79 132 304 275 Qwen-2.5-3B Phi-mini-3.8B Both: Deterministic 8.08 6.41 128 224 200 Qwen-2.5-3B Phi-mini-3.8B Both: Exploratory 7.96 6.29 136 339 303 Qwen-2.5-3B Phi-mini-3.8B Both: Det. & Exp. 8.13 6.46 133 241 216 Qwen-2.5-3B Phi-mini-3.8B Both: Exp. & Det. 8.83 7.16 134 307 276 Qwen-2.5-1.5B Qwen-2.5-3B Both: Default 21.79 2.04 154 573 537 Qwen-2.5-1.5B Qwen-2.5-3B Both: Deterministic 21.92 2.17 180 500 471 Qwen-2.5-1.5B Qwen-2.5-3B Both: Exploratory 21.79 2.04 165 680 639 Qwen-2.5-1.5B Qwen-2.5-3B Both: Det. & Exp. 20.58 0.83 174 525 483 Qwen-2.5-1.5B Qwen-2.5-3B Both: Exp. & Det. 22.25 2.50 189 640 608 Llama-3.1-3B Llama-3.1-8B Both: Default 11.08 1.13 264 422 381 Llama-3.1-3B Llama-3.1-8B Both: Deterministic 11.41 1.46 278 348 316 Llama-3.1-3B Llama-3.1-8B Both: Exploratory 11.50 1.55 241 490 424 Llama-3.1-3B Llama-3.1-8B Both: Det. & Exp. 11.54 1.59 259 420 362 Llama-3.1-3B Llama-3.1-8B Both: Exp. & Det. 11.00 1.05 298 411 364 Qwen-2.5-7B Qwen-2.5-14B Both: Default 7.26 4.09 88 165 159 Qwen-2.5-7B Qwen-2.5-14B Both: Deterministic 6.92 3.75 83 119 112 Qwen-2.5-7B Qwen-2.5-14B Both: Exploratory 6.46 3.29 111 168 153 Qwen-2.5-7B Qwen-2.5-14B Both: Det. & Exp. 7.50 4.33 92 155 148 Qwen-2.5-7B Qwen-2.5-14B Both: Exp. & Det. 7.71 4.54 78 143 133
Agent 1 Agent 2 Agent 3 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 2400) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Default 0.21 583 473 299 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Deterministic 4.42 0 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Exploratory 4.04 645 578 344 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 1 Det. & 2 Exp. 2.12 667 467 296 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 2 Det. & 1 Exp. 0.63 454 394 279 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Default 11.04 446 995 676 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Deterministic 5.29 0 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Exploratory 11.33 357 1159 774 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 1 Det. & 2 Exp. 11.67 394 1116 756 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 2 Det. & 1 Exp. 11.17 265 793 514 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Default 5.63 273 423 326 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Deterministic 5.63 0 0 0 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Exploratory 6.25 223 537 404 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 1 Det. & 2 Exp. 6.79 235 428 343 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 2 Det. & 1 Exp. 5.37 202 274 208 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Default 7.17 84 272 209 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Deterministic 5.00 0 0 0 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Exploratory 6.34 117 270 220 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 1 Det. & 2 Exp. 6.63 120 267 214 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 2 Det. & 1 Exp. 5.80 97 181 135 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Default 6.13 55 166 140 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Deterministic 4.75 0 3 1 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Exploratory 5.50 69 188 159 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 1 Det. & 2 Exp. 5.42 72 172 143 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 2 Det. & 1 Exp. 5.42 48 105 81 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Default 1.00 48 112 96 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Deterministic 0.33 0 0 0 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Exploratory 1.00 46 123 109 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 1 Det. & 2 Exp. 1.42 42 131 106 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 2 Det. & 1 Exp. 0.66 40 75 60 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Default 6.79 226 389 284 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Deterministic 5.75 0 3 1 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Exploratory 6.83 219 423 327 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 1 Det. & 2 Exp. 6.41 232 390 293 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 2 Det. & 1 Exp. 6.12 191 292 202 Mistral-7B Mistral-7B Mistral-7B Default 8.99 312 979 525 Mistral-7B Mistral-7B Mistral-7B Deterministic 0.67 0 8 3 Mistral-7B Mistral-7B Mistral-7B Exploratory 12.00 325 1234 628 Mistral-7B Mistral-7B Mistral-7B 1 Det. & 2 Exp. 8.88 349 1046 544 Mistral-7B Mistral-7B Mistral-7B 2 Det. & 1 Exp. 7.96 232 855 458 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Default 5.87 454 733 476 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Deterministic 5.00 0 0 0 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Exploratory 5.04 520 857 544 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 1 Det. & 2 Exp. 4.50 515 744 493 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 2 Det. & 1 Exp. 5.66 309 493 322 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Default 7.05 345 572 407 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Deterministic 5.42 0 0 0 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Exploratory 5.46 385 624 446 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 1 Det. & 2 Exp. 6.50 374 588 413 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 2 Det. & 1 Exp. 5.50 335 414 269
Agent 1 Agent 2 Agent 3 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 2400) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-3B Default 1.75 338 1356 951 Qwen-2.5-0.5B Qwen-2.5-1.5B Llama-3.1-3B Default 1.79 596 1086 718 Qwen-2.5-0.5B Qwen-2.5-1.5B Phi-mini-3.8B Default 6.80 503 1168 857 Qwen-2.5-0.5B Qwen-2.5-3B Llama-3.1-3B Default 2.13 364 1202 895 Qwen-2.5-0.5B Qwen-2.5-3B Phi-mini-3.8B Default 1.83 353 1190 946 Qwen-2.5-0.5B Llama-3.1-3B Phi-mini-3.8B Default 6.50 536 980 724 Qwen-2.5-1.5B Qwen-2.5-3B Llama-3.1-3B Default 2.25 321 1048 773 Qwen-2.5-1.5B Qwen-2.5-3B Phi-mini-3.8B Default 3.83 299 857 692 Qwen-2.5-1.5B Llama-3.1-3B Phi-mini-3.8B Default 0.08 405 1079 766 Qwen-2.5-3B Phi-mini-3.8B Llama-3.1-3B Default 5.66 255 825 653 Qwen-2.5-3B Qwen-2.5-3B Phi-mini-3.8B Default 7.04 291 461 340 Qwen-2.5-3B Phi-mini-3.8B Phi-mini-3.8B Default 5.79 279 424 317 Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Default 7.88 456 1197 794 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-1.5B Default 4.79 589 969 607
Agent 1 Agent 2 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 2376) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Default 52.90 1.73 550 482 Qwen-2.5-0.5B Qwen-2.5-0.5B Deterministic 53.24 1.39 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Exploratory 49.07 5.56 628 530 Qwen-2.5-0.5B Qwen-2.5-0.5B Det. & Exp. 52.99 1.64 572 516 Qwen-2.5-1.5B Qwen-2.5-1.5B Default 86.15 0.47 415 403 Qwen-2.5-1.5B Qwen-2.5-1.5B Deterministic 84.60 2.02 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Exploratory 83.42 3.20 574 547 Qwen-2.5-1.5B Qwen-2.5-1.5B Det. & Exp. 86.62 449 434 Qwen-2.5-3B Qwen-2.5-3B Default 94.02 0.96 117 114 Qwen-2.5-3B Qwen-2.5-3B Deterministic 93.35 0.29 0 0 Qwen-2.5-3B Qwen-2.5-3B Exploratory 94.15 1.09 158 157 Qwen-2.5-3B Qwen-2.5-3B Det. & Exp. 94.07 1.01 126 124 Qwen-2.5-7B Qwen-2.5-7B Default 96.17 1.48 39 37 Qwen-2.5-7B Qwen-2.5-7B Deterministic 96.55 1.86 0 0 Qwen-2.5-7B Qwen-2.5-7B Exploratory 96.93 2.24 57 53 Qwen-2.5-7B Qwen-2.5-7B Det. & Exp. 96.46 1.77 35 34 Qwen-2.5-14B Qwen-2.5-14B Default 98.19 2.53 21 21 Qwen-2.5-14B Qwen-2.5-14B Deterministic 97.77 2.11 0 0 Qwen-2.5-14B Qwen-2.5-14B Exploratory 98.15 2.49 20 20 Qwen-2.5-14B Qwen-2.5-14B Det. & Exp. 97.94 2.28 24 24 Qwen-2.5-32B Qwen-2.5-32B Default 98.53 0.21 14 13 Qwen-2.5-32B Qwen-2.5-32B Deterministic 98.36 0.04 0 0 Qwen-2.5-32B Qwen-2.5-32B Exploratory 98.53 0.21 14 14 Qwen-2.5-32B Qwen-2.5-32B Det. & Exp. 98.36 0.04 10 8 Phi-mini-3.8B Phi-mini-3.8B Default 95.88 3.92 71 60 Phi-mini-3.8B Phi-mini-3.8B Deterministic 95.37 3.41 0 0 Phi-mini-3.8B Phi-mini-3.8B Exploratory 94.74 2.78 126 116 Phi-mini-3.8B Phi-mini-3.8B Det. & Exp. 94.95 2.99 89 78 Mistral-7B Mistral-7B Default 81.06 0.04 227 219 Mistral-7B Mistral-7B Deterministic 80.43 0.59 0 0 Mistral-7B Mistral-7B Exploratory 80.18 0.84 261 251 Mistral-7B Mistral-7B Det. & Exp. 82.41 1.39 240 235 Llama-3.1-3B Llama-3.1-3B Default 87.71 3.07 163 153 Llama-3.1-3B Llama-3.1-3B Deterministic 86.66 2.02 0 0 Llama-3.1-3B Llama-3.1-3B Exploratory 88.09 3.45 216 208 Llama-3.1-3B Llama-3.1-3B Det. & Exp. 86.91 2.27 181 172 Llama-3.1-8B Llama-3.1-8B Default 94.44 5.34 79 75 Llama-3.1-8B Llama-3.1-8B Deterministic 93.64 4.54 0 0 Llama-3.1-8B Llama-3.1-8B Exploratory 93.60 4.50 118 109 Llama-3.1-8B Llama-3.1-8B Det. & Exp. 94.53 5.43 95 93
Agent 1 Agent 2 Agent Settings Accuracy Lower Upper Debate Sycophancy CI IC Debate Rounds (Avg / 2376) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Default 22.35 9.64 804 760 Qwen-2.5-0.5B Qwen-2.5-1.5B Deterministic 24.75 7.24 734 711 Qwen-2.5-0.5B Qwen-2.5-1.5B Exploratory 18.56 13.43 899 828 Qwen-2.5-0.5B Qwen-2.5-1.5B Det. & Exp. 20.58 11.41 846 790 Qwen-2.5-0.5B Qwen-2.5-1.5B Exp. & Det. 23.02 8.97 829 794 Qwen-2.5-1.5B Llama-3.1-3B Default 1.93 3.91 376 357 Qwen-2.5-1.5B Llama-3.1-3B Deterministic 1.51 3.49 242 239 Qwen-2.5-1.5B Llama-3.1-3B Exploratory 1.43 3.41 483 457 Qwen-2.5-1.5B Llama-3.1-3B Det. & Exp. 0.37 1.35 290 277 Qwen-2.5-1.5B Llama-3.1-3B Exp. & Det. 1.09 2.07 447 433 Qwen-2.5-3B Phi-mini-3.8B Default 2.18 3.28 135 132 Qwen-2.5-3B Phi-mini-3.8B Deterministic 1.85 2.95 106 102 Qwen-2.5-3B Phi-mini-3.8B Exploratory 2.18 3.28 184 178 Qwen-2.5-3B Phi-mini-3.8B Det. & Exp. 1.85 2.95 148 148 Qwen-2.5-3B Phi-mini-3.8B Exp. & Det. 2.69 3.79 146 139 Qwen-2.5-1.5B Qwen-2.5-3B Default 5.26 1.18 363 359 Qwen-2.5-1.5B Qwen-2.5-3B Deterministic 5.97 0.47 263 254 Qwen-2.5-1.5B Qwen-2.5-3B Exploratory 5.17 1.27 498 487 Qwen-2.5-1.5B Qwen-2.5-3B Det. & Exp. 6.14 0.20 294 286 Qwen-2.5-1.5B Qwen-2.5-3B Exp. & Det. 5.89 0.45 469 466 Llama-3.1-3B Llama-3.1-8B Default 7.15 2.69 184 179 Llama-3.1-3B Llama-3.1-8B Deterministic 6.48 2.02 138 133 Llama-3.1-3B Llama-3.1-8B Exploratory 5.97 1.51 202 192 Llama-3.1-3B Llama-3.1-8B Det. & Exp. 6.35 1.89 152 149 Llama-3.1-3B Llama-3.1-8B Exp. & Det. 7.32 2.86 229 222 Qwen-2.5-7B Qwen-2.5-14B Default 3.25 2.28 55 55 Qwen-2.5-7B Qwen-2.5-14B Deterministic 2.95 1.98 48 47 Qwen-2.5-7B Qwen-2.5-14B Exploratory 2.70 1.73 67 66 Qwen-2.5-7B Qwen-2.5-14B Det. & Exp. 2.74 1.77 49 48 Qwen-2.5-7B Qwen-2.5-14B Exp. & Det. 2.78 1.81 49 48
Agent 1 Agent 2 Agent Settings MAD Accuracy Debate Sycophancy CI IC Debate (RCR Prompting) Rounds (Avg / 2376) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Default 51.30 3.33 990 642 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Deterministic 53.24 1.39 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Exploratory 46.80 7.83 1099 655 Qwen-2.5-0.5B Qwen-2.5-0.5B 1 Det. & 2 Exp. 48.99 5.64 1053 658 Qwen-2.5-0.5B Qwen-2.5-0.5B 2 Det. & 1 Exp. 50.80 3.83 793 495 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Default 87.37 0.75 717 573 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Deterministic 84.60 2.02 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Exploratory 85.61 1.01 1011 795 Qwen-2.5-1.5B Qwen-2.5-1.5B 1 Det. & 2 Exp. 86.32 0.30 834 672 Qwen-2.5-1.5B Qwen-2.5-1.5B 2 Det. & 1 Exp. 86.53 0.09 587 451 Qwen-2.5-3B Qwen-2.5-3B Both: Default 94.87 1.81 196 165 Qwen-2.5-3B Qwen-2.5-3B Both: Deterministic 93.35 0.29 0 0 Qwen-2.5-3B Qwen-2.5-3B Both: Exploratory 94.28 1.22 252 206 Qwen-2.5-3B Qwen-2.5-3B 1 Det. & 2 Exp. 94.70 1.64 238 195 Qwen-2.5-3B Qwen-2.5-3B 2 Det. & 1 Exp. 93.94 0.88 162 117 Qwen-2.5-7B Qwen-2.5-7B Both: Default 96.21 1.52 69 58 Qwen-2.5-7B Qwen-2.5-7B Both: Deterministic 96.17 1.48 0 0 Qwen-2.5-7B Qwen-2.5-7B Both: Exploratory 96.55 1.86 86 71 Qwen-2.5-7B Qwen-2.5-7B 1 Det. & 2 Exp. 96.55 1.86 78 65 Qwen-2.5-7B Qwen-2.5-7B 2 Det. & 1 Exp. 96.34 1.65 56 40 Qwen-2.5-14B Qwen-2.5-14B Both: Default 98.15 2.49 29 26 Qwen-2.5-14B Qwen-2.5-14B Both: Deterministic 97.77 2.11 0 0 Qwen-2.5-14B Qwen-2.5-14B Both: Exploratory 98.19 2.53 40 36 Qwen-2.5-14B Qwen-2.5-14B 1 Det. & 2 Exp. 98.02 2.36 40 31 Qwen-2.5-14B Qwen-2.5-14B 2 Det. & 1 Exp. 97.81 2.15 28 25 Qwen-2.5-32B Qwen-2.5-32B Both: Default 98.57 0.25 15 13 Qwen-2.5-32B Qwen-2.5-32B Both: Deterministic 98.36 0.04 0 0 Qwen-2.5-32B Qwen-2.5-32B Both: Exploratory 98.48 0.16 14 14 Qwen-2.5-32B Qwen-2.5-32B 1 Det. & 2 Exp. 98.48 0.16 15 12 Qwen-2.5-32B Qwen-2.5-32B 2 Det. & 1 Exp. 98.32 0.00 9 6 Phi-mini-3.8B Phi-mini-3.8B Both: Default 95.79 3.83 138 105 Phi-mini-3.8B Phi-mini-3.8B Both: Deterministic 95.37 3.41 0 0 Phi-mini-3.8B Phi-mini-3.8B Both: Exploratory 94.91 2.95 234 185 Phi-mini-3.8B Phi-mini-3.8B 1 Det. & 2 Exp. 96.34 4.38 189 149 Phi-mini-3.8B Phi-mini-3.8B 2 Det. & 1 Exp. 95.92 3.96 115 83 Llama-3.1-3B Llama-3.1-3B Both: Default 87.33 2.69 292 227 Llama-3.1-3B Llama-3.1-3B Both: Deterministic 87.63 2.99 0 0 Llama-3.1-3B Llama-3.1-3B Both: Exploratory 87.71 3.07 415 323 Llama-3.1-3B Llama-3.1-3B 1 Det. & 2 Exp. 87.58 2.94 328 259 Llama-3.1-3B Llama-3.1-3B 2 Det. & 1 Exp. 88.47 3.83 236 169 Llama-3.1-8B Llama-3.1-8B Both: Default 93.86 4.76 139 102 Llama-3.1-8B Llama-3.1-8B Both: Deterministic 93.64 4.54 0 0 Llama-3.1-8B Llama-3.1-8B Both: Exploratory 94.19 5.09 190 141 Llama-3.1-8B Llama-3.1-8B 1 Det. & 2 Exp. 94.11 5.01 185 143 Llama-3.1-8B Llama-3.1-8B 2 Det. & 1 Exp. 94.49 5.39 139 89 Mistral-7B Mistral-7B Both: Default 82.20 1.18 469 342 Mistral-7B Mistral-7B Both: Deterministic 80.43 0.59 0 0 Mistral-7B Mistral-7B Both: Exploratory 82.66 1.64 566 429 Mistral-7B Mistral-7B 1 Det. & 2 Exp. 82.37 1.35 506 376 Mistral-7B Mistral-7B 2 Det. & 1 Exp. 81.69 0.67 346 230
Agent 1 Agent 2 Agent 3 MAD Accuracy Debate Sycophancy CI IC Debate (RCR Prompting) Rounds (Avg / 2376) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-3B 0.34 145 1377 1153 Qwen-2.5-0.5B Qwen-2.5-1.5B Llama-3.1-3B 387 1223 1006 Qwen-2.5-0.5B Qwen-2.5-1.5B Phi-mini-3.8B 0.97 184 1379 1156 Qwen-2.5-0.5B Qwen-2.5-3B Llama-3.1-3B 1.86 213 1221 1070 Qwen-2.5-0.5B Qwen-2.5-3B Phi-mini-3.8B 3.58 299 1157 1024 Qwen-2.5-0.5B Llama-3.1-3B Phi-mini-3.8B 0.17 238 559 479 Qwen-2.5-1.5B Qwen-2.5-3B Llama-3.1-3B 1.22 189 560 479 Qwen-2.5-1.5B Qwen-2.5-3B Phi-mini-3.8B 2.48 103 509 449 Qwen-2.5-1.5B Llama-3.1-3B Phi-mini-3.8B 0.17 238 559 479 Qwen-2.5-3B Phi-mini-3.8B Llama-3.1-3B 1.01 162 332 283 Qwen-2.5-3B Qwen-2.5-3B Phi-mini-3.8B 2.82 86 253 214 Qwen-2.5-3B Phi-mini-3.8B Phi-mini-3.8B 3.28 71 227 180 Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 2.00 300 1229 1012 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-1.5B 14.19 616 1400 982
Agent 1 Agent 2 Agent Settings MAD Accuracy Debate Sycophancy CI IC Debate (ARC-Challenge) Rounds (Avg / 1172) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Default 39.51 1.54 265 228 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Deterministic 40.78 2.81 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Exploratory 37.54 0.43 309 245 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Det. & Exp. 39.85 1.88 259 227 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Default 70.90 1.69 249 242 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Deterministic 67.58 1.63 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Exploratory 68.52 0.69 296 275 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Det. & Exp. 69.88 0.67 262 252 Qwen-2.5-3B Qwen-2.5-3B Both: Default 85.41 1.88 114 111 Qwen-2.5-3B Qwen-2.5-3B Both: Deterministic 84.13 0.60 0 0 Qwen-2.5-3B Qwen-2.5-3B Both: Exploratory 84.64 1.11 116 109 Qwen-2.5-3B Qwen-2.5-3B Both: Det. & Exp. 83.70 0.17 79 73 Qwen-2.5-7B Qwen-2.5-7B Both: Default 91.55 4.33 46 45 Qwen-2.5-7B Qwen-2.5-7B Both: Deterministic 91.21 3.99 0 0 Qwen-2.5-7B Qwen-2.5-7B Both: Exploratory 91.64 4.42 53 51 Qwen-2.5-7B Qwen-2.5-7B Both: Det. & Exp. 92.06 4.84 48 43 Qwen-2.5-14B Qwen-2.5-14B Both: Default 94.54 4.27 24 24 Qwen-2.5-14B Qwen-2.5-14B Both: Deterministic 94.37 4.10 0 0 Qwen-2.5-14B Qwen-2.5-14B Both: Exploratory 93.77 3.50 24 24 Qwen-2.5-14B Qwen-2.5-14B Both: Det. & Exp. 94.71 4.44 22 21 Qwen-2.5-32B Qwen-2.5-32B Both: Default 98.53 3.25 14 13 Qwen-2.5-32B Qwen-2.5-32B Both: Deterministic 98.36 3.08 0 0 Qwen-2.5-32B Qwen-2.5-32B Both: Exploratory 98.53 3.25 14 14 Qwen-2.5-32B Qwen-2.5-32B Both: Det. & Exp. 98.36 3.08 10 8 Phi-mini-3.8B Phi-mini-3.8B Both: Default 90.10 5.37 75 66 Phi-mini-3.8B Phi-mini-3.8B Both: Deterministic 88.91 4.18 0 0 Phi-mini-3.8B Phi-mini-3.8B Both: Exploratory 87.03 2.30 107 100 Phi-mini-3.8B Phi-mini-3.8B Both: Det. & Exp. 88.05 3.32 69 62 Llama-3.1-3B Llama-3.1-3B Both: Default 75.77 2.65 130 126 Llama-3.1-3B Llama-3.1-3B Both: Deterministic 74.66 1.54 0 0 Llama-3.1-3B Llama-3.1-3B Both: Exploratory 76.19 3.07 166 149 Llama-3.1-3B Llama-3.1-3B Both: Det. & Exp. 75.60 2.48 129 124 Llama-3.1-8B Llama-3.1-8B Both: Default 87.20 9.55 91 88 Llama-3.1-8B Llama-3.1-8B Both: Deterministic 85.75 8.10 0 0 Llama-3.1-8B Llama-3.1-8B Both: Exploratory 85.07 7.42 96 94 Llama-3.1-8B Llama-3.1-8B Both: Det. & Exp. 86.86 9.21 84 80 Mistral-7B Mistral-7B Both: Default 70.48 1.71 145 137 Mistral-7B Mistral-7B Both: Deterministic 68.26 0.51 0 0 Mistral-7B Mistral-7B Both: Exploratory 72.78 4.01 185 177 Mistral-7B Mistral-7B Both: Det. & Exp. 70.82 2.05 151 142
Agent 1 Agent 2 Agent Settings MAD Accuracy Debate Sycophancy CI IC Debate (ARC-Challenge) Rounds (Avg / 1172) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Default 58.28 20.31 10.93 193 401 369 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Deterministic 63.57 25.60 5.64 169 375 357 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Exploratory 55.80 17.83 13.41 211 418 368 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Det. & Exp. 60.32 22.35 8.89 181 397 360 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Exp. & Det. 61.43 23.46 7.78 197 409 387 Qwen-2.5-1.5B Llama-3.1-3B Both: Default 72.35 3.14 0.77 143 216 207 Qwen-2.5-1.5B Llama-3.1-3B Both: Deterministic 74.91 5.70 1.79 135 191 185 Qwen-2.5-1.5B Llama-3.1-3B Both: Exploratory 73.12 3.91 153 281 265 Qwen-2.5-1.5B Llama-3.1-3B Both: Det. & Exp. 76.02 6.81 2.90 127 219 205 Qwen-2.5-1.5B Llama-3.1-3B Both: Exp. & Det. 74.15 4.94 1.03 135 291 274 Qwen-2.5-3B Phi-mini-3.8B Both: Default 87.97 4.44 3.24 59 133 130 Qwen-2.5-3B Phi-mini-3.8B Both: Deterministic 88.57 5.04 3.84 58 110 107 Qwen-2.5-3B Phi-mini-3.8B Both: Exploratory 87.03 3.50 2.30 72 173 160 Qwen-2.5-3B Phi-mini-3.8B Both: Det. & Exp. 87.80 4.27 3.07 59 141 139 Qwen-2.5-3B Phi-mini-3.8B Both: Exp. & Det. 89.85 6.32 5.12 50 143 137 Qwen-2.5-1.5B Qwen-2.5-3B Both: Default 82.25 13.04 1.28 80 247 243 Qwen-2.5-1.5B Qwen-2.5-3B Both: Deterministic 82.59 13.38 0.94 80 205 200 Qwen-2.5-1.5B Qwen-2.5-3B Both: Exploratory 81.91 12.70 1.62 94 317 310 Qwen-2.5-1.5B Qwen-2.5-3B Both: Det. & Exp. 83.45 14.24 0.08 66 227 219 Qwen-2.5-1.5B Qwen-2.5-3B Both: Exp. & Det. 83.62 14.41 0.09 67 328 320 Llama-3.1-3B Llama-3.1-8B Both: Default 81.66 8.54 4.01 114 141 133 Llama-3.1-3B Llama-3.1-8B Both: Deterministic 80.46 7.34 2.81 120 135 124 Llama-3.1-3B Llama-3.1-8B Both: Exploratory 75.68 2.56 1.97 107 160 151 Llama-3.1-3B Llama-3.1-8B Both: Det. & Exp. 80.12 7.00 2.47 117 138 132 Llama-3.1-3B Llama-3.1-8B Both: Exp. & Det. 80.97 7.85 3.32 109 159 154 Qwen-2.5-7B Qwen-2.5-14B Both: Default 93.43 6.21 3.16 35 54 53 Qwen-2.5-7B Qwen-2.5-14B Both: Deterministic 93.60 6.38 3.33 24 59 58 Qwen-2.5-7B Qwen-2.5-14B Both: Exploratory 94.45 7.23 4.18 27 67 65 Qwen-2.5-7B Qwen-2.5-14B Both: Det. & Exp. 93.00 5.78 2.73 37 50 49 Qwen-2.5-7B Qwen-2.5-14B Both: Exp. & Det. 93.77 6.55 3.50 26 58 58
Agent 1 Agent 2 Agent 3 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 1172) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Default 2.82 535 484 283 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Deterministic 2.81 0 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Exploratory 2.65 587 528 303 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 1 Det. & 2 Exp. 0.77 523 503 306 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 2 Det. & 1 Exp. 0.26 404 353 219 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Default 3.32 206 454 343 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Deterministic 1.63 0 0 0 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Exploratory 2.89 235 611 433 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 1 Det. & 2 Exp. 2.72 229 520 386 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 2 Det. & 1 Exp. 1.61 163 345 245 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Default 2.22 79 197 156 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Deterministic 0.60 0 0 0 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Exploratory 2.73 96 229 167 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 1 Det. & 2 Exp. 2.73 106 193 149 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 2 Det. & 1 Exp. 1.20 71 131 101 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Default 4.59 56 84 66 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Deterministic 3.39 0 0 0 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Exploratory 4.50 66 85 65 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 1 Det. & 2 Exp. 3.82 60 80 68 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 2 Det. & 1 Exp. 4.08 40 57 40 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Default 3.93 27 54 45 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Deterministic 4.10 0 0 0 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Exploratory 4.53 28 50 39 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 1 Det. & 2 Exp. 4.27 22 41 33 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 2 Det. & 1 Exp. 4.44 10 32 26 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Default 0.54 22 36 28 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Deterministic 0.45 0 0 0 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Exploratory 0.28 28 35 32 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 1 Det. & 2 Exp. 0.28 30 29 25 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 2 Det. & 1 Exp. 0.71 13 18 14 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Default 4.18 69 130 104 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Deterministic 4.18 0 0 0 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Exploratory 4.01 85 196 151 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 1 Det. & 2 Exp. 4.01 74 155 121 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 2 Det. & 1 Exp. 4.35 54 109 81 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Default 2.65 177 244 190 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Deterministic 1.71 0 0 0 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Exploratory 2.39 196 303 210 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 1 Det. & 2 Exp. 2.05 223 262 192 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 2 Det. & 1 Exp. 2.14 118 162 117 Mistral-7B Mistral-7B Mistral-7B Default 1.96 213 292 207 Mistral-7B Mistral-7B Mistral-7B Deterministic 0.51 0 0 0 Mistral-7B Mistral-7B Mistral-7B Exploratory 2.90 232 360 249 Mistral-7B Mistral-7B Mistral-7B 1 Det. & 2 Exp. 2.48 209 317 227 Mistral-7B Mistral-7B Mistral-7B 2 Det. & 1 Exp. 1.71 142 214 136 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Default 9.81 98 145 107 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Deterministic 8.78 0 0 0 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Exploratory 8.36 127 187 150 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 1 Det. & 2 Exp. 9.04 114 174 128 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 2 Det. & 1 Exp. 8.02 115 119 73
Agent 1 Agent 2 Agent 3 Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 1172) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-3B 0.94 148 820 629 Qwen-2.5-0.5B Qwen-2.5-1.5B Llama-3.1-3B 5.12 311 641 489 Qwen-2.5-0.5B Qwen-2.5-1.5B Phi-mini-3.8B 1.97 170 804 621 Qwen-2.5-0.5B Qwen-2.5-3B Llama-3.1-3B 3.84 208 699 581 Qwen-2.5-0.5B Qwen-2.5-3B Phi-mini-3.8B 2.22 133 722 631 Qwen-2.5-0.5B Llama-3.1-3B Phi-mini-3.8B 6.32 238 683 559 Qwen-2.5-1.5B Qwen-2.5-3B Llama-3.1-3B 1.19 180 447 358 Qwen-2.5-1.5B Qwen-2.5-3B Phi-mini-3.8B 2.64 105 423 358 Qwen-2.5-1.5B Llama-3.1-3B Phi-mini-3.8B 3.00 195 412 341 Qwen-2.5-3B Phi-mini-3.8B Llama-3.1-3B 2.14 143 319 244 Qwen-2.5-3B Qwen-2.5-3B Phi-mini-3.8B 3.15 110 225 170 Qwen-2.5-3B Phi-mini-3.8B Phi-mini-3.8B 4.60 81 214 174 Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 0.59 231 686 523 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-1.5B 13.24 393 680 451
Agent 1 Agent 2 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 1221) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Default 39.80 3.31 306 240 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Deterministic 40.87 4.38 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Exploratory 33.50 2.99 338 257 Qwen-2.5-0.5B Qwen-2.5-0.5B Both: Det. & Exp. 41.93 5.44 355 289 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Default 67.40 0.88 154 154 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Deterministic 68.14 1.62 2 1 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Exploratory 67.24 0.72 217 201 Qwen-2.5-1.5B Qwen-2.5-1.5B Both: Det. & Exp. 66.67 0.15 166 158 Qwen-2.5-3B Qwen-2.5-3B Both: Default 74.37 1.71 128 123 Qwen-2.5-3B Qwen-2.5-3B Both: Deterministic 74.77 2.11 0 0 Qwen-2.5-3B Qwen-2.5-3B Both: Exploratory 73.87 1.21 127 120 Qwen-2.5-3B Qwen-2.5-3B Both: Det. & Exp. 75.51 2.85 127 123 Qwen-2.5-7B Qwen-2.5-7B Both: Default 81.57 2.01 66 64 Qwen-2.5-7B Qwen-2.5-7B Both: Deterministic 81.65 2.09 0 0 Qwen-2.5-7B Qwen-2.5-7B Both: Exploratory 81.90 2.34 78 75 Qwen-2.5-7B Qwen-2.5-7B Both: Det. & Exp. 82.56 3.00 62 61 Qwen-2.5-14B Qwen-2.5-14B Both: Default 83.37 1.00 43 41 Qwen-2.5-14B Qwen-2.5-14B Both: Deterministic 83.70 1.33 0 0 Qwen-2.5-14B Qwen-2.5-14B Both: Exploratory 83.21 0.84 66 62 Qwen-2.5-14B Qwen-2.5-14B Both: Det. & Exp. 83.87 1.50 59 54 Qwen-2.5-32B Qwen-2.5-32B Both: Default 86.24 0.48 47 46 Qwen-2.5-32B Qwen-2.5-32B Both: Deterministic 85.75 0.01 0 0 Qwen-2.5-32B Qwen-2.5-32B Both: Exploratory 86.24 0.48 46 43 Qwen-2.5-32B Qwen-2.5-32B Both: Det. & Exp. 86.57 0.81 55 46 Phi-mini-3.8B Phi-mini-3.8B Both: Default 71.66 1.78 100 79 Phi-mini-3.8B Phi-mini-3.8B Both: Deterministic 72.24 2.36 0 0 Phi-mini-3.8B Phi-mini-3.8B Both: Exploratory 73.87 3.99 141 121 Phi-mini-3.8B Phi-mini-3.8B Both: Det. & Exp. 73.22 3.34 124 105 Llama-3.1-3B Llama-3.1-3B Both: Default 68.55 3.51 117 110 Llama-3.1-3B Llama-3.1-3B Both: Deterministic 67.40 2.36 0 0 Llama-3.1-3B Llama-3.1-3B Both: Exploratory 66.75 1.71 131 122 Llama-3.1-3B Llama-3.1-3B Both: Det. & Exp. 67.73 2.69 113 109 Mistral-7B Mistral-7B Both: Default 66.34 1.79 64 57 Mistral-7B Mistral-7B Both: Deterministic 66.99 2.44 0 0 Mistral-7B Mistral-7B Both: Exploratory 65.11 0.56 85 80 Mistral-7B Mistral-7B Both: Det. & Exp. 66.42 1.87 89 81 Llama-3.1-8B Llama-3.1-8B Both: Default 74.28 1.26 114 106 Llama-3.1-8B Llama-3.1-8B Both: Deterministic 75.43 2.41 2 1 Llama-3.1-8B Llama-3.1-8B Both: Exploratory 74.86 1.84 139 130 Llama-3.1-8B Llama-3.1-8B Both: Det. & Exp. 74.45 1.43 112 102
Agent 1 Agent 2 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 1221) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Default 56.92 20.43 9.60 370 345 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Deterministic 58.39 21.90 8.13 326 295 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Exploratory 56.91 20.42 9.61 430 377 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Det. & Exp. 57.08 20.59 9.44 371 332 Qwen-2.5-0.5B Qwen-2.5-1.5B Both: Exp. & Det. 57.49 21.00 9.03 407 379 Qwen-2.5-1.5B Llama-3.1-3B Both: Default 66.83 0.31 1.79 170 165 Qwen-2.5-1.5B Llama-3.1-3B Both: Deterministic 68.63 2.11 3.59 198 184 Qwen-2.5-1.5B Llama-3.1-3B Both: Exploratory 67.08 0.56 2.04 237 223 Qwen-2.5-1.5B Llama-3.1-3B Both: Det. & Exp. 69.78 3.26 4.74 203 193 Qwen-2.5-1.5B Llama-3.1-3B Both: Exp. & Det. 67.73 1.21 2.69 219 200 Qwen-2.5-3B Phi-mini-3.8B Both: Default 75.02 2.36 5.14 158 150 Qwen-2.5-3B Phi-mini-3.8B Both: Deterministic 76.09 3.43 6.21 161 154 Qwen-2.5-3B Phi-mini-3.8B Both: Exploratory 74.69 2.03 4.81 177 167 Qwen-2.5-3B Phi-mini-3.8B Both: Det. & Exp. 75.76 3.10 5.88 191 179 Qwen-2.5-3B Phi-mini-3.8B Both: Exp. & Det. 75.10 2.44 5.22 162 156 Qwen-2.5-1.5B Qwen-2.5-3B Both: Default 73.87 7.35 1.21 225 217 Qwen-2.5-1.5B Qwen-2.5-3B Both: Deterministic 74.94 8.42 2.28 191 187 Qwen-2.5-1.5B Qwen-2.5-3B Both: Exploratory 74.12 7.60 1.46 279 264 Qwen-2.5-1.5B Qwen-2.5-3B Both: Det. & Exp. 74.04 7.52 1.38 208 204 Qwen-2.5-1.5B Qwen-2.5-3B Both: Exp. & Det. 74.94 8.42 2.28 251 246 Llama-3.1-3B Llama-3.1-8B Both: Default 72.24 7.20 0.78 165 153 Llama-3.1-3B Llama-3.1-8B Both: Deterministic 73.79 8.75 0.77 190 183 Llama-3.1-3B Llama-3.1-8B Both: Exploratory 72.15 7.11 0.87 167 157 Llama-3.1-3B Llama-3.1-8B Both: Det. & Exp. 70.68 5.64 2.34 162 154 Llama-3.1-3B Llama-3.1-8B Both: Exp. & Det. 73.96 8.92 0.94 200 193 Qwen-2.5-7B Qwen-2.5-14B Both: Default 83.37 3.81 1.00 98 96 Qwen-2.5-7B Qwen-2.5-14B Both: Deterministic 83.78 4.22 1.41 101 95 Qwen-2.5-7B Qwen-2.5-14B Both: Exploratory 84.19 4.63 1.82 112 110 Qwen-2.5-7B Qwen-2.5-14B Both: Det. & Exp. 83.37 3.81 1.00 103 99 Qwen-2.5-7B Qwen-2.5-14B Both: Exp. & Det. 83.29 3.73 0.92 95 93
Agent 1 Agent 2 Agent 3 Agent Settings Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 1221) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Default 1.27 545 538 327 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Deterministic 3.31 0 0 0 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B Exploratory 3.89 580 604 336 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 1 Det. & 2 Exp. 0.28 569 558 317 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-0.5B 2 Det. & 1 Exp. 1.02 433 420 237 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Default 2.28 193 333 264 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Deterministic 1.38 0 3 1 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B Exploratory 1.05 256 429 315 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 1 Det. & 2 Exp. 2.03 211 346 270 Qwen-2.5-1.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 2 Det. & 1 Exp. 2.03 172 244 179 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Default 2.52 147 225 180 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Deterministic 1.62 0 0 0 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B Exploratory 1.71 164 248 196 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 1 Det. & 2 Exp. 2.36 166 211 163 Qwen-2.5-3B Qwen-2.5-3B Qwen-2.5-3B 2 Det. & 1 Exp. 3.10 116 163 115 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Default 2.34 85 122 96 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Deterministic 2.01 0 0 0 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B Exploratory 2.42 99 147 117 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 1 Det. & 2 Exp. 1.85 98 124 99 Qwen-2.5-7B Qwen-2.5-7B Qwen-2.5-7B 2 Det. & 1 Exp. 2.18 84 89 65 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Default 0.68 84 85 69 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Deterministic 1.50 0 0 0 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B Exploratory 0.76 76 100 75 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 1 Det. & 2 Exp. 1.17 74 93 77 Qwen-2.5-14B Qwen-2.5-14B Qwen-2.5-14B 2 Det. & 1 Exp. 1.58 45 56 46 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Default 0.24 61 80 67 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Deterministic 0.01 0 0 0 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B Exploratory 0.81 56 87 74 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 1 Det. & 2 Exp. 0.24 61 71 57 Qwen-2.5-32B Qwen-2.5-32B Qwen-2.5-32B 2 Det. & 1 Exp. 0.32 35 50 41 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Default 3.34 170 171 121 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Deterministic 3.83 0 0 0 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B Exploratory 4.08 161 231 170 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 1 Det. & 2 Exp. 5.30 134 217 159 Phi-mini-3.8B Phi-mini-3.8B Phi-mini-3.8B 2 Det. & 1 Exp. 3.83 107 137 97 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Default 3.35 210 237 169 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Deterministic 3.02 0 0 0 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B Exploratory 2.61 250 261 190 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 1 Det. & 2 Exp. 2.04 213 225 165 Llama-3.1-3B Llama-3.1-3B Llama-3.1-3B 2 Det. & 1 Exp. 2.69 132 149 105 Mistral-7B Mistral-7B Mistral-7B Default 2.28 121 137 99 Mistral-7B Mistral-7B Mistral-7B Deterministic 2.20 0 0 0 Mistral-7B Mistral-7B Mistral-7B Exploratory 1.05 179 167 119 Mistral-7B Mistral-7B Mistral-7B 1 Det. & 2 Exp. 0.89 157 144 97 Mistral-7B Mistral-7B Mistral-7B 2 Det. & 1 Exp. 2.20 81 98 68 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Default 2.90 147 211 148 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Deterministic 2.82 0 9 3 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B Exploratory 1.10 203 246 168 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 1 Det. & 2 Exp. 2.49 173 233 161 Llama-3.1-8B Llama-3.1-8B Llama-3.1-8B 2 Det. & 1 Exp. 2.49 118 150 92
Agent 1 Agent 2 Agent 3 Accuracy Debate Sycophancy CI IC Debate Rounds (Avg / 1221) Helped (Avg) (Overall) Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-3B 35.99 228 748 563 Qwen-2.5-0.5B Qwen-2.5-1.5B Llama-3.1-3B 28.54 343 622 480 Qwen-2.5-0.5B Qwen-2.5-1.5B Phi-mini-3.8B 34.11 246 691 537 Qwen-2.5-0.5B Qwen-2.5-3B Llama-3.1-3B 36.07 234 697 544 Qwen-2.5-0.5B Qwen-2.5-3B Phi-mini-3.8B 35.66 243 629 517 Qwen-2.5-0.5B Llama-3.1-3B Phi-mini-3.8B 32.63 298 617 483 Qwen-2.5-1.5B Qwen-2.5-3B Llama-3.1-3B 6.86 230 399 305 Qwen-2.5-1.5B Qwen-2.5-3B Phi-mini-3.8B 9.16 202 382 303 Qwen-2.5-1.5B Llama-3.1-3B Phi-mini-3.8B 4.57 260 347 273 Qwen-2.5-3B Phi-mini-3.8B Llama-3.1-3B 1.54 222 334 253 Qwen-2.5-3B Qwen-2.5-3B Phi-mini-3.8B 2.11 200 256 193 Qwen-2.5-3B Phi-mini-3.8B Phi-mini-3.8B 3.43 183 258 186 Qwen-2.5-0.5B Qwen-2.5-1.5B Qwen-2.5-1.5B 28.37 267 576 447 Qwen-2.5-0.5B Qwen-2.5-0.5B Qwen-2.5-1.5B 18.63 384 651 438
Appendix F Additional Results
F.1 Original MAD Results
We also report our experiments with the original Multi-Agent Debate (MAD) framework across various model sizes and architectures. Table 33 presents the results on three challenging reasoning benchmarks: GSM-Plus, GSM8K, and ARC-Challenge.
F.2 Majority Vote@3 Results
To further investigate the impact of stochastic diversity on model performance, we report results on a Majority Vote@3 approach where we sample three independent responses from each model and take a majority vote to determine the final answer. Table 34 presents these results across five benchmarks: GSM8K, GSM-Plus, ARC-Easy, ARC-Challenge, and CommonsenseQA.
The results demonstrate that simple ensemble-based approaches can significantly boost performance without requiring multi-agent debate or model fine-tuning. Across all model sizes and architectures, Majority Vote@3 consistently outperforms single-sample inference. The relative improvements are most pronounced for smaller models, with Qwen-2.5-0.5B gaining up to 4.27 percentage points on ARC-Challenge and Qwen-2.5-1.5B showing similar substantial improvements across benchmarks.
Interestingly, this pattern holds across model families. Llama-3.1-3B, Phi-3.5-mini, and Mistral-7B all exhibit significant gains when using majority voting, suggesting that the benefits of ensemble diversity transcend specific model architectures. The results also indicate diminishing returns for larger models—Qwen-2.5-14B shows more modest improvements compared to its smaller counterparts, likely because these larger models already produce more consistent answers across samples.
These findings highlight an important baseline for our research: simple ensemble methods provide strong performance improvements with minimal computational overhead during inference. However, they still require multiple forward passes for each query, motivating our DTE approach that aims to distill these benefits into a single model through training on debate traces.
F.3 Scaling Results for Multiple Agents
We investigated how performance scales with increasing numbers of debating agents (1-7) across different model sizes and reasoning benchmarks. Table 35 presents these results, revealing several important trends in multi-agent scaling behavior.
First, we observe that performance generally improves as we add more agents to the debate, but with diminishing returns. The most significant gains occur when moving from a single agent (equivalent to standard inference) to two agents, with more modest improvements as additional agents join the debate. For example, on GSM8K, Qwen-2.5-1.5B shows a substantial jump from 62.77% (1 agent) to 71.57% (2 agents), but only incremental improvements thereafter.
Second, the benefits of additional agents vary across tasks. On more complex tasks like GSM-Plus, we see continued performance improvements even with 7 agents, particularly for larger models. Qwen-2.5-14B shows its peak GSM-Plus performance with 7 agents (78.08%), suggesting that more difficult problems benefit from extended multi-agent collaboration. In contrast, on simpler tasks like ARC-Easy, performance plateaus more quickly.
Third, we find that model size influences scaling behavior. Smaller models like Qwen-2.5-1.5B show more variability in performance as agents are added, with occasional performance drops when moving from 3 to 4 agents. Larger models exhibit more stable scaling patterns, suggesting that they can more consistently integrate insights from multiple debate participants.
These results have important implications for our DTE framework. They demonstrate that while adding more agents generally improves performance, the computational costs may outweigh the benefits beyond 3-5 agents for most applications. This insight helped inform our design choices in balancing performance gains against computational efficiency in our final framework.
| Model Configuration | Debate Performance Metrics | ||||||||
| Agent 1 | Agent 2 | Debate Setting | Accuracy | Delta | Debate Rounds | Sycophancy | Correct→Incorrect | Incorrect→Correct | Net Benefit |
| GSM-Plus | |||||||||
| Qwen-2.5-0.5B | Qwen-2.5-0.5B | exploratory | 28.12% | 3.33 | 3.48 | 6906 | 261 | 575 | 432 |
| Qwen-2.5-1.5B | Qwen-2.5-1.5B | exploratory | 46.50% | 4.50 | 2.33 | 5642 | 194 | 861 | 670 |
| Qwen-2.5-3B | Qwen-2.5-3B | exploratory | 66.79% | 5.04 | 1.34 | 5315 | 231 | 373 | 187 |
| Qwen-2.5-7B | Qwen-2.5-7B | exploratory | 69.71% | 1.09 | 0.76 | 2967 | 102 | 200 | 121 |
| Qwen-2.5-14B | Qwen-2.5-14B | exploratory | 76.92% | 5.13 | 0.61 | 2722 | 119 | 151 | 47 |
| Phi-mini-3.8B | Phi-mini-3.8B | exploratory | 65.79% | 2.37 | 1.07 | 3620 | 180 | 272 | 136 |
| Llama-3.1-3B | Llama-3.1-3B | exploratory | 42.42% | 3.25 | 2.07 | 5507 | 379 | 369 | 238 |
| Mistral-7B | Mistral-7B | exploratory | 26.35% | 11.31 | 1.85 | 4500 | 210 | 290 | 115 |
| Llama-3.1-8B | Llama-3.1-8B | exploratory | 57.63% | 2.01 | 1.75 | 5667 | 273 | 585 | 351 |
| GSM8K | |||||||||
| Qwen-2.5-0.5B | Qwen-2.5-0.5B | exploratory | 45.56% | 3.56 | 2.85 | 3469 | 175 | 427 | 328 |
| Qwen-2.5-1.5B | Qwen-2.5-1.5B | exploratory | 65.81% | 3.04 | 1.99 | 3471 | 144 | 650 | 489 |
| Qwen-2.5-3B | Qwen-2.5-3B | exploratory | 86.96% | 1.82 | 0.63 | 1390 | 82 | 165 | 97 |
| Qwen-2.5-7B | Qwen-2.5-7B | exploratory | 91.74% | 1.07 | 0.38 | 930 | 64 | 93 | 33 |
| Qwen-2.5-14B | Qwen-2.5-14B | exploratory | 94.39% | 1.59 | 0.18 | 448 | 30 | 48 | 18 |
| Phi-mini-3.8B | Phi-mini-3.8B | exploratory | 88.17% | 1.29 | 0.45 | 1050 | 65 | 120 | 65 |
| Llama-3.1-3B | Llama-3.1-3B | exploratory | 67.63% | 4.92 | 1.51 | 2418 | 238 | 215 | 127 |
| Mistral-7B | Mistral-7B | exploratory | 43.44% | 22.06 | 1.65 | 2100 | 175 | 235 | 95 |
| Llama-3.1-8B | Llama-3.1-8B | exploratory | 83.02% | 1.29 | 0.94 | 1587 | 93 | 308 | 236 |
| ARC-Challenge | |||||||||
| Qwen-2.5-0.5B | Qwen-2.5-0.5B | exploratory | 38.65% | 0.68 | 1.88 | 2728 | 272 | 308 | 232 |
| Qwen-2.5-1.5B | Qwen-2.5-1.5B | exploratory | 74.15% | 0.94 | 0.85 | 1671 | 121 | 231 | 156 |
| Qwen-2.5-3B | Qwen-2.5-3B | exploratory | 85.41% | 1.88 | 0.57 | 1227 | 94 | 135 | 57 |
| Qwen-2.5-7B | Qwen-2.5-7B | exploratory | 91.47% | 6.25 | 0.23 | 501 | 41 | 49 | 13 |
| Qwen-2.5-14B | Qwen-2.5-14B | exploratory | 94.54% | 4.27 | 0.15 | 326 | 31 | 37 | 9 |
| Phi-mini-3.8B | Phi-mini-3.8B | exploratory | 87.46% | 2.73 | 0.15 | 313 | 24 | 47 | 25 |
| Llama-3.1-3B | Llama-3.1-3B | exploratory | 76.37% | 3.25 | 0.73 | 1525 | 111 | 155 | 69 |
| Mistral-7B | Mistral-7B | exploratory | 73.29% | 4.52 | 0.40 | 795 | 63 | 114 | 73 |
| Llama-3.1-8B | Llama-3.1-8B | exploratory | 86.09% | 8.44 | 0.27 | 514 | 31 | 84 | 58 |
| Model | Accuracy (%) on Benchmarks | ||||
| GSM8K | GSM-Plus | ARC-E | ARC-C | CQA | |
| Qwen-2.5-0.5B | 49.73 | 30.54 | 58.71 | 42.92 | 42.51 |
| Qwen-2.5-1.5B | 75.82 | 52.08 | 87.12 | 73.55 | 69.62 |
| Qwen-2.5-3B | 86.28 | 64.08 | 94.19 | 84.13 | 76.90 |
| Qwen-2.5-7B | 92.19 | 70.46 | 96.46 | 91.21 | 82.88 |
| Qwen-2.5-14B | 94.09 | 72.54 | 98.44 | 94.20 | 82.15 |
| Llama-3.1-3B | 77.03 | 52.79 | 88.51 | 75.00 | 69.94 |
| Llama-3.1-8B | 85.82 | 60.88 | 93.56 | 83.11 | 74.86 |
| Phi-3.5-mini | 87.87 | 65.79 | 96.00 | 86.95 | 75.10 |
| Mistral-7B | 56.86 | 36.88 | 87.58 | 75.68 | 69.04 |
| Model | Number of Agents | ||||||
| 1 | 2 | 3 | 4 | 5 | 6 | 7 | |
| GSM8K Accuracy (%) | |||||||
| Qwen-2.5-1.5B | 62.77 | 71.57 | 75.13 | 75.89 | 75.13 | 74.98 | 76.50 |
| Qwen-2.5-3B | 85.14 | 85.52 | 87.64 | 87.11 | 87.04 | 86.66 | 87.11 |
| Qwen-2.5-7B | 90.67 | 91.21 | 92.42 | 92.49 | 92.57 | 92.34 | 92.72 |
| Qwen-2.5-14B | 92.80 | 93.33 | 94.84 | 94.31 | 94.69 | 94.62 | 94.24 |
| GSM-Plus Accuracy (%) | |||||||
| Qwen-2.5-1.5B | 42.00 | 51.62 | 53.33 | 50.62 | 54.21 | 51.50 | 52.67 |
| Qwen-2.5-3B | 61.75 | 67.79 | 68.00 | 64.21 | 69.71 | 64.88 | 68.54 |
| Qwen-2.5-7B | 68.62 | 74.17 | 74.96 | 70.88 | 71.08 | 71.38 | 76.00 |
| Qwen-2.5-14B | 71.79 | 77.25 | 72.29 | 72.83 | 73.29 | 73.38 | 78.08 |
| ARC-Challenge Accuracy (%) | |||||||
| Qwen-2.5-1.5B | 69.21 | 68.52 | 72.10 | 71.50 | 72.53 | 71.50 | 72.10 |
| Qwen-2.5-3B | 82.53 | 84.64 | 86.26 | 85.75 | 86.26 | 86.95 | 87.03 |
| Qwen-2.5-7B | 87.22 | 91.64 | 91.72 | 91.47 | 92.06 | 91.38 | 92.32 |
| Qwen-2.5-14B | 90.27 | 93.77 | 94.80 | 95.14 | 94.20 | 94.62 | 94.28 |
| ARC-Easy Accuracy (%) | |||||||
| Qwen-2.5-1.5B | 86.62 | 83.42 | 85.61 | 86.32 | 87.46 | 86.57 | 87.16 |
| Qwen-2.5-3B | 93.06 | 94.15 | 94.28 | 94.32 | 94.82 | 94.91 | 94.99 |
| Qwen-2.5-7B | 94.69 | 96.93 | 96.55 | 96.34 | 96.42 | 96.25 | 96.59 |
| Qwen-2.5-14B | 95.66 | 98.15 | 98.19 | 98.23 | 98.15 | 98.19 | 98.23 |